### v1

# DiffoRA: Enabling Parameter-Efficient LLM Fine-Tuning via Differential Low-Rank Matrix Adaptation

Abstract
The Parameter-Efficient Fine-Tuning (PEFT) methods have been extensively researched for large language models in the downstream tasks. Among all the existing approaches, the Low-Rank Adaptation (LoRA) has gained popularity for its streamlined design by incorporating low-rank matrices into existing pre-trained models. Though effective, LoRA allocates every module an identical low-rank matrix, which ignores the varying properties and contributions across different components. Moreover, the existing adaptive LoRA solutions rely highly on intuitive importance scoring indicators to adjust the interior rank of the decomposition matrices. In this paper, we propose a new PEFT scheme called DiffoRA, which is theoretically grounded and enables module-wise adoption of LoRA. At the core of our DiffoRA lies a Differential Adaptation Matrix (DAM) to determine which module is the most suitable and essential for fine-tuning. We explain how the designed matrix impacts the convergence rate and generalization capability of a pre-trained model. Furthermore, we construct the DAM via continuous relaxation and discretization with weight-sharing optimizations. We fully implement our DiffoRA and design comprehensive experiments to evaluate its performance. The experimental results demonstrate that our approach achieves the best model accuracy over all the state-of-the-art baselines across various benchmarks.
1 Introduction
In recent years, large language models (LLMs) have gained significant attraction across various domains of natural language processing, demonstrating remarkable capabilities in tasks such as text generation [18, 25], machine translation [33], sentiment analysis [32], and question-answering [40]. Due to the large model size, fine-tuning an LLM to some downstream tasks will invoke a large number of parameters, e.g., up to 175 billion parameters for GPT-3 in a fully fine-tuning. As a result, Parameter-Efficient Fine-Tuning (PEFT) becomes increasingly paramount due to its alignment with the LLM demands.
A plethora of PEFT methods have been proposed, which can be categorized into two classes. Some approaches slightly modify the model structure by adding small trainable modules and keeping the rest of the models unchanged [19, 16]. Another line of work aims to efficiently capture the incremental parameter updates without modifying the model structure [7]. Among all the existing work, the Low-Rank Adaptation [12] is widely acknowledged due to its effectiveness and satisfying performance. Unlike the previous work, LoRA incurs some low-rank decomposition matrices to parameterize the incremental updates and realizes comparable results with 70% less overhead than the full fine-tuning. This innovative approach paves the way for more effective utilization of LLMs in practice.
Though effective, there are some limitations of LoRA-based methods. The most fundamental one is that LoRA treats all the modules in the network equally by adding the decomposition matrices to all the trainable modules in the network. Consequently, it ignores the varying properties and contributions of different modules in fine-tuning. Some works have been proposed to adaptively utilize LoRA. Generally speaking, these approaches managed to score the importance of different modules and adjust the interior rank of each decomposition matrix accordingly [37, 36, 21]. For example, Zhang et al. [36] designed AdaLoRA, which utilizes the singular value of the low-rank matrices to adjust the rank. These methods rely highly on the intuitive metrics (e.g., singular values or norms [23, 38]) to score the importance across different modules, which overlook theoretical relationships between the evaluation metrics and the incremental updates of the model. Moreover, the performance of the existing methods can be further enhanced.
Our Objectives. To alleviate the above limitations, in this paper, we aim to answer the following key question: Can we construct a theoretically grounded PEFT method to enable adaptively adoption of the low-rank decomposition matrices, thus we can only fine-tune the modules that are most necessary and essential?
Our contributions. To this end, in this paper, we shed some light on improving the PEFT performance based on LoRA and propose a novel method called DiffoRA. We argue that instead of adjusting the interior rank of every decomposition matrix like previous works, it suffices to adopt the low-rank matrices module-wisely. Thus, at the core of our approach lies a Differentiable Adaptation Matrix (DAM) that determines the importance of each module, so that which one needs to be fine-tuned and vice versa. Unlike the previous intuitive metrics, DAM is processed to be differentiable w.r.t. the incremental updates of the low-rank matrices, thus enabling the capture of the essential characteristics of the module importance. Specifically, we first analyze how the DAM impacts the model performance (i.e., the convergence rate and generalization capabilities) during fine-tuning in theory. We then elaborate on the algorithm to approximate the NP-hard problem by constructing the DAM via continuous relaxation and discretization. To alleviate the possible discrepancy problem and enhance the robustness of DiffoRA, we further incorporate the weight-sharing strategy as optimization. We fully implement our DiffoRA and design extensive experiments to evaluate its performance on two widely adopted benchmarks and multiple tasks. The experimental results show that our scheme is consistently better than all the existing baselines. For instance, DiffoRA achieves 0.81% higher model accuracy on CoLA task [31] than the state-of-the-art method. In all, the contributions of this paper are as follows.
  • We propose DiffoRA, a novel PEFT method for LLMs that is adaptive and theoretically grounded. Our approach is built atop a newly designed differentiable matrix (i.e., DAM), which enables adaptive adoption of the low-rank decomposition matrices.
  • We theoretically explain how the DAM impacts the performance of a LoRA-based fine-tuning model in terms of the convergence rate and generalization capability.
  • We fully implement our DiffoRA and evaluate its performance on two widely adopted benchmarks that contain multiple tasks. The experimental results demonstrate that DiffoRA works consistently better than all the baselines.
2 Related work
2.1 LLM and fine-tuning
LLM has captured considerable public interest due to its success in multiple regions. The PEFT is essential for LLMs due to the huge amount of parameters. Some previous works have been proposed to fine-tune the LLMs using specifically designed modules that are added to LLMs. The fine-tuning of LLMs is thus converted to the adjustments of these small modules. For instance, multiple methods [28, 19, 16] insert dataset-dependent small modules or vectors between the layers to decrease the parameter amounts. Another line of work models the incremental updates of the fine-tuning procedure to make it parameter-efficient. Guo et al. [7] propose to use a task-specific difference vector to extend and fine-tune the base model. There are also some methods that fine-tune parts of the parameters [6], e.g., the bias of FFN [35] or the last quarter [15].
2.2 Low-rank adaptation and optimizations
To further reduce the computational and storage cost, Hu et al. [12] proposed LoRA, in which they designed a low-rank decomposition matrix to model the incremental updates during fine-tuning. A plethora of work has been proposed to optimize LoRA and reduce the parameter amount [14, 8, 34, 2, 29]. One of the limitations of LoRA is that it treats all the modules equally, which omits the variations of the modules in LLMs. To address this issue, a few works have been proposed to realize an adaptive LoRA. Without loss of generality, these methods first evaluate the importance of the modules and then adjust the interior rank of the decomposition matrices accordingly. For instance, AdaLoRA [36] utilizes the singular value to score the importance, while some other approaches adopt the Frobenius norm [23] or norm of the outputs [38] as the metrics. These indicators are intuitively adopted and only utilize partial information of the incremental updates. There are also a few works that utilize the training procedures to determine the module importance [3, 37, 21], yet they still focus on modifying the interior ranks of the decomposition matrices.
3 Theoretical analysis
3.1 Preliminaries
Notations. We define $[n]=\{1,2,\ldots,n\}$. We denote the vectors and matrices as the lower and uppercase bold font, respectively. For instance, $\bm{x}$ is a vector with entry $x_{i}$, and $\bm{M}$ is a matrix with entry $[\bm{M}]_{ij}$. The minimum eigenvalue of $\bm{M}$ is denoted as $\lambda_{\min}(\bm{M})$. $\|\cdot\|_{2}$ is used to represent the $l_{2}$ norm of a vector. $N(\bm{0},\bm{I})$ and $U\{S\}$ represent the standard Gaussian distribution and uniform distribution over a set $S$, respectively. We denote by $\bm{X}=\{(\bm{x}_{i},y_{i})|\bm{x}_{i}\in\mathbb{R}^{d\times 1},y_{i}\in\mathbb{R},i\in[n]\}$ the training set, where $\bm{x}_{i}$ and $y_{i}$ represent the $i$-th data and label. $\mathbb{I}\{\cdot\}$ represents the indicator function that demonstrates the event occurrence, such that for event $\mathcal{A}$, $\mathbb{I}\{\mathcal{A}\}=1$ if and only if $\mathcal{A}$ happened, otherwise it equals to 0. $P(\mathcal{A})$ represents the probability of $\mathcal{A}$ occurred event.
Neural networks and gram matrix. For input $\bm{x}\in\mathbb{R}^{d\times 1}$, weight vector $\bm{w}\in\mathbb{R}^{d\times 1}$ in the weight matrix $\bm{W}\in\mathbb{R}^{d\times m}$, and output weight $\bm{a}\in\mathbb{R}^{m\times 1}$, we denote $f(\bm{W},\bm{a},\bm{x})$ as a neural network with a single hidden layer such that
$f(\bm{W},\bm{a},\bm{x})=\frac{1}{\sqrt{m}}\sum\nolimits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}^{T}\bm{x})$ (1)
where $\sigma$ is the activation function. In this paper, we primarily consider the ReLU function, which is one of the most adopted activation functions in the literature, i.e., $\sigma(z)=z\mathbb{I}\{z>0\}$. Given a training set $\bm{X}$, the optimization goal is to minimize the empirical risk loss function
$\mathcal{L}(\bm{W},\bm{a})=\sum\nolimits_{i=1}^{n}\frac{1}{2}(f(\bm{W},\bm{a},\bm{x}_{i})-y_{i})^{2}$ (2)
In this work, we aim to construct a module-wise DAM $\bm{\Gamma}\in\mathbb{R}^{L\times N}$ with entries $\gamma_{i,j}\in\{0,1\}$ to determine the necessity of each module for fine-tuning in LLM, where $L$ is the number of layers and $N$ is the module amount in each layer. To further analyze the relationship between the model performance and the model with $\bm{\Gamma}$, we expand the definitions in Eq. 1 and Eq. 2 such that
$f(\bm{W},\bm{a},\bm{x};\bm{\Gamma},\bm{W}_{0})=\frac{1}{\sqrt{m}}\sum_{r=1}^{m}a_{r}\sigma((\bm{w}_{0}+\bm{\Gamma}\bm{w}_{r})^{T}\bm{x})$ (3)
$\mathcal{L}(\bm{W},\bm{a};\bm{\Gamma},\bm{W}_{0})=\sum\nolimits_{i=1}^{n}\frac{1}{2}(f(\bm{W},\bm{a},\bm{x}_{i};\bm{\Gamma},\bm{W}_{0})-y_{i})^{2}$
where $\bm{w}_{i}$ is the $i$-th row of $\bm{W}$. Furthermore, we follow the definitions in [4] and define the matrices $\mathbf{H}^{\infty}_{\Gamma,w_{0}}$ and $\mathbf{H}^{\infty}_{w_{0}}$ based on $\bm{\Gamma}$ such that
Definition 1 (Gram Matrix).
For a neural network with a single hidden layer, the gram matrix $\mathbf{H}^{\infty}_{\Gamma,w_{0}}\in\mathbb{R}^{n\times n}$ induced by the ReLU activation function on a training set $\bm{X}:=\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$ with entry
$[\bm{H}_{\Gamma,w_{0}}^{\infty}]_{ij}=$ $\mathbb{E}_{\bm{w}\sim N(\bm{0},\bm{I})}[\bm{x}_{i}^{T}\bm{x}_{j}\cdot$ (4)
$\mathbb{I}\{(\bm{w}_{0}+\bm{\Gamma}\bm{w})^{T}\bm{x}_{i}\geq 0,(\bm{w}_{0}+\bm{\Gamma}\bm{w})^{T}\bm{x}_{j}\geq 0\}]$
$=$ $\bm{x}_{i}^{T}\bm{x}_{j}[\bm{I}^{\bm{\Gamma}\bm{w}}]_{ij}$
We further construct $\mathbf{H}_{w_{0}}^{\infty}$ with entry $[\mathbf{H}_{w_{0}}^{\infty}]_{ij}$ such that
$[\bm{H}^{\infty}_{w_{0}}]_{ij}=$ $\mathbb{E}_{\bm{w}\sim N(\bm{0},\bm{I})}[\bm{x}_{i}^{T}\bm{x}_{j}\cdot$ (5)
$\mathbb{I}\{(\bm{w}_{0}+\bm{w})^{T}\bm{x}_{i}\geq 0,(\bm{w}_{0}+\bm{w})^{T}\bm{x}_{j}\geq 0\}]$
$=$ $\bm{x}_{i}^{T}\bm{x}_{j}[\bm{I}^{\bm{w}}]_{ij}$
where $\bm{I}^{\bm{\Gamma w}}$ and $\bm{I}^{\bm{w}}$ are the expectations of the indicator matrices corresponding to vectors $\bm{\Gamma w}$ and $\bm{w}$, respectively. We denote $\lambda_{0}:=\lambda_{\min}(\mathbf{H}_{w_{0}}^{\infty})$, and $\lambda_{0}^{\Gamma}:=\lambda_{\min}(\mathbf{H}_{\Gamma,w_{0}}^{\infty})$.
Recall LoRA. LoRA [12] utilizes two matrices $\bm{A}\in\mathbb{R}^{r\times k},\bm{B}\in\mathbb{R}^{d\times r}$ to substitute the parameters’ increments. For $\bm{h}=\bm{W}^{(0)}\bm{x}$, the forward pass in LoRA is
$\bm{h}=\bm{W}^{0}\bm{x}+\Delta\bm{x}=\bm{W}^{0}\bm{x}+\bm{B}\cdot\bm{A}\bm{x}$ (6)
where $d,k\ll r$, $\bm{A}$ is usually initialized by following Gaussian distribution and $\bm{B}$ is initialized with zeros. LoRA adopts this modification equally to all the modules in the model. In the following, we denote $\Delta\bm{W}$ as $\bm{B}\cdot\bm{A}$.
3.2 Main theorems
Technical intuitions. In this section, we focus on analyzing the model performance regarding two aspects, i.e., the convergence rate and the generalization capability. The existing schemes that adaptively adopt LoRA focus on adjusting the interior rank of the matrices based on some intuitive evaluation metrics. In contrast to these methods, we argue that adopting LoRA module-wisely in each layer is sufficient. Thus, our intuition in this work is to exert a binary matrix (i.e., $\bm{\Gamma}$) on the model structure to realize a “selective” fine-tuning of the model. In this section, we first assume there exists an appropriate algorithm to construct $\bm{\Gamma}$ and explain why adopting it can lead to better model performance.
Our analysis is established on the intuitive observations that fine-tuning a pre-trained model can be viewed as training a well-initialized model. We construct the theories to explain how the $\bm{\Gamma}$ of a well-initialized over-parameterized Neural Network (NN) impacts the convergence rate and generalization capability. In the following section, we will first present the theorem regarding $\lambda_{0}$ and the model performance then demonstrate our main theorems.
Theoretical results. The over-parameterized NNs are widely adopted and competitive in hierarchical feature extraction due to the large number of parameters they contain. We use the architecture with wide hidden layers in this section to establish our theories of the matrix $\mathbf{\Gamma}$. This network is one of the most fundamental structures of the over-parameterized NN and is proved to be tractable in training [17, 13]. Based on this insight, we present the following theorem [4] about the convergence rate of the NN with a single hidden layer as follows.
Theorem 1.
If gram matrix $\bm{H}^{\infty}\succ 0$, $\|\bm{x}_{i}\|_{2}=1$, $|y_{i}|<C$ for some constant $C$ and $i\in[n]$, hidden nodes $m=\Omega\left(\frac{n^{6}}{\lambda_{\min}(\bm{H}^{\infty})^{4}\delta^{3}}\right)$, and i.i.d. initialize $\bm{w}_{r}\sim N(\bm{0},\bm{I})$, $a_{r}\sim U\{[-1,1]\}$ for $r\in[m]$, then with probability at least $1-\delta$ over the initialization, the following inequality holds:
$\|f(\bm{W}(t),\bm{a},\bm{X})-\bm{y}\|_{2}^{2}$ (7)
$\leq$ $\exp{(-\lambda_{\min}(\bm{H}^{\infty})t)}\|f(\bm{W}(0),\bm{a},\bm{X})-\bm{y}\|_{2}^{2}$
where
$\bm{H}^{\infty}:=\mathbb{E}_{\bm{w}\sim N(\bm{0},\bm{I})}[\bm{x}_{i}^{T}\bm{x}_{j}\mathbb{I}\{(\bm{w}^{T}\bm{x}_{i}\geq 0,\bm{w}^{T}\bm{x}_{j}\geq 0\}].$
The inequality in the above theorem demonstrates that the minimum eigenvalue of the Gram matrix positively affects the training convergence rate of the network. Thus, it is viable for us to evaluate the convergence rate of a network from the minimum eigenvalue of the Gram matrix.
Based on Theorem 1, we analyze the relationship of the minimum eigenvalue between the adaptive matrix (i.e., $\lambda^{\Gamma}_{0}$) and the original matrix (i.e., $\lambda_{0}$). We summarize our results in the following theorem:
Theorem 2.
Suppose $f$ is an NN with a single hidden layer and ReLU activation function. Assume $\bm{X}\in\mathbb{R}^{d\times n}$, $\bm{w}(0)\sim N(\bm{0},\bm{I})$, hidden nodes $m=\Omega\left(\frac{n^{6}d^{2}}{(\lambda_{0}^{\Gamma})^{4}\delta^{3}}\right)$, and $\bm{I}^{\bm{\Gamma w}}-\bm{I}^{\bm{w}}\succeq 0$, then the following formula holds with probability at least $1-\delta$ over the initialization
$\|f(\bm{W}(t),\bm{a},\bm{X};\bm{\Gamma},\bm{W}_{0})-\bm{y}\|_{2}^{2}$ (8)
$\leq$ $\exp(-\lambda_{0}^{\Gamma}t)\|f(\bm{W}(0),\bm{a},\bm{X};\bm{\Gamma},\bm{W}_{0})-\bm{y}\|_{2}^{2}$
where $\lambda_{0}^{\Gamma}\geq\lambda_{0}$.
Proof sketch. The key of the proof is to find the relationship between $\lambda_{0}$ and $\lambda_{0}^{\Gamma}$ by Weyl inequalities [10]. We provide the full proof in Sec. A.1.
Assume that we can establish a matrix that satisfies this requirement, then Theorem 2 demonstrates that the minimum eigenvalue of the Gram matrix of the network with $\mathbf{\Gamma}$ is larger than the $\lambda_{0}$ without the selective matrix, thus leading to a higher convergence rate.
Other than the convergence rate, we also analyze the relationship between $\lambda_{0}^{\Gamma}$ and the generalization capability of the over-parameterized NN. We present the results in the following theorem:
Theorem 3.
For an over-parameterized neural network with the loss on the testing set as $\mathcal{L}(\bm{W},\bm{a};\bm{\Gamma},\bm{W}_{0})$. Let $\bm{y}=(y_{1},...,y_{N})^{T}$, and $\eta=\kappa C_{1}\sqrt{\bm{y}^{T}(\bm{H}_{\Gamma,w_{0}}^{\infty})^{-1}\bm{y}}/(m\sqrt{N})$ for some small enough absolute constant $\kappa$, where $\eta$ denotes the step of SGD. Under the assumption of Theorem 2, for any $\delta\in(0,e^{-1}]$, there exists $m^{\ast}(\delta,N,\lambda_{0}^{\Gamma})$, such that if $m\geq m^{*}$, then with probability at least $1-\delta$, we have
$\mathbb{E}[L(\bm{W},\bm{a};\bm{\Gamma},\bm{W}_{0})]\leq$ $\mathcal{O}(C^{\prime}\sqrt{\frac{\bm{y}^{T}\bm{y}}{\lambda_{0}^{\Gamma}N}})+\mathcal{O}(\sqrt{\frac{\log(1/\delta)}{N}})$ (9)
where $\lambda_{0}^{\Gamma}>\lambda_{0}$, $C,C^{\prime}$, and $\delta$ are constants.
Proof sketch. The proof of the above theorem derives from Corollary 3.10 of [1] and Section D.2 of [39]. We present the detailed proof in Sec. A.2.
Similarly to Theorem 2, the above theorem indicates that the adoption of $\bm{\Gamma}$ can enhance the generalization capability of the model, as shown in the last inequality.
Overall, in this section, we have proven that a selective matrix $\bm{\Gamma}$ will result in a higher convergence rate and better generalization capability of the network, thus leading to enhanced model performance. The main theorems in Theorem 2 and Theorem 3 theoretically grounded the effectiveness of the selective matrix. The question remains is how to construct $\bm{\Gamma}$ for the network efficiently. We will provide solutions and detailed algorithms in the next section.
4 Design of DiffoRA
We now present the concrete method to construct the module-wise selective matrix $\mathbf{\Gamma}$. The proposed approach is called DiffoRA, which is built atop the DAM (i.e., Differentiable Adaptation Matrix). To capture the information of the incremental updates modeled by the low-rank decomposition matrices, DiffoRA views the elements $\gamma$ in the module-wise DAM $\bm{\Gamma}$ as trainable parameters. More concretely, as shown in Fig. 1, DiffoRA approximates the NP-hard problem by invoking the following two stages: (i) Relaxation and optimization, which aims to map $\bm{\Gamma}$ to a continuous space (i.e., $\bar{\bm{\Gamma}}$) so that it is differentiable and can be further optimized; and (ii) Discretization and fine-tuning, which binarizes $\bar{\bm{\Gamma}}$ to determine the most essential module for fine-tuning. We will first elaborate on each stage respectively, and then introduce the weight-sharing optimization.
4.1 Continuous relaxation
Algorithm 1 DiffoRA
0:  Pre-trained model with $L$ layers $\mathcal{M}^{L}$; Candidate fine-tuning modules $M$, where $|M|=L\times N$, $N$ is the number of candidate modules in each layers; Training dataset and valid dataset $\bm{X}_{train}$ and $\bm{X}_{valid}$; Training epochs and valid epochs, $T$ and $V$; The learning rate $\eta$; Sample rate $\rho$; LoRA rank: $r_{l}$; Share rank: $r_{s}$.
1:  // Stage 1: Continuous Relaxation
2:  Create the hyperparameters $\bar{\bm{\Gamma}}\in\mathbb{R}^{L\times N}$.
3:  for $v=1;v<V;v++$ do
4:     Update hyperparameters $\bar{\bm{\Gamma}}$ as $\bar{\bm{\Gamma}}_{v}=\bar{\bm{\Gamma}}_{v}-\eta\nabla_{\bar{\bm{\Gamma}}}\mathcal{L}_{valid}(\bm{X}_{valid};\Delta\bm{W},\bar{\bm{\Gamma}}_{v-1})$
5:     for $t=1;t<T;t++$ do
6:        Update low-rank matrix $\Delta\bm{W}$ as follow:
$\Delta\bm{W}=\Delta\bm{W}-\eta\nabla_{\Delta\bm{W}}\mathcal{L}_{train}(\bm{X}_{train};\Delta\bm{W},\bar{\bm{\Gamma}}_{v})$
7:     end for
8:  end for
9:  // Stage 2: Discretization and Fine-Tuning
10:  Select the Top-K modules of each layer in $\mathcal{M}^{L}$ according to the $\bar{\bm{\Gamma}}$, where $K=\lfloor\rho\cdot N\rfloor$.
11:  Add a low-rank matrix $\Delta\bm{W}=\bm{BA}$ to the selected module and add the weight sharing matrix $\Delta\bm{W}_{s}=\bm{B}_{s}\bm{A}_{s}$ to the remaining modules, where $\bm{B}^{T},\bm{A}\in\mathbb{R}^{d\times r_{l}}$, and $\bm{B}^{T}_{s},\bm{A}_{s}\in\mathbb{R}^{d\times r_{s}}$.
12:  for $t=1;t<T;t++$ do
13:     Update low-rank matrix $\Delta\bm{W}$ as follow:
$\Delta\bm{W}=\Delta\bm{W}-\eta\nabla_{\Delta\bm{W}}\mathcal{L}_{train}(\bm{X}_{train};\Delta\bm{W})$
14:  end for
15:  return  Fine-tuned model $\mathcal{M}^{L}$.
Figure 1: The framework of DiffoRA, contains two stages. In stage one (left part), the initialized adaptive matrix is continuously relaxed, i.e., C.R. in the figure. The DAM is then differentiable and can be updated. In stage two (right part), the obtained DAM is discretized to binary. All the modules corresponding to entry one will be fine-tuned using decomposition matrices.
The first stage of DiffoRA is illustrated in the left part of Fig. 1. In contrast with the existing work, we point out that it is unnecessary to allocate the rank for every module. Instead, we can construct a selective matrix called DAM module-wisely that determines which module needs to be fine-tuned and which need not. The final output DAM $\bm{\Gamma}$ should be a binary matrix, in which the “ones” (resp. “zeros”) indicate that the corresponding entries will (resp. will not) be fine-tuned in the following procedures. However, directly generating a binary matrix is non-trivial and lacks foundation, i.e., it is an NP-hard problem. To this end, in our design of DiffoRA, we first relax the range of the elements in $\bm{\Gamma}\in\{0,1\}^{L\times N}$ to $\bm{\bar{\Gamma}}\in[0,1]^{L\times N}$ continuously and view all the row vectors in $\bar{\bm{\Gamma}}$ as the hyperparameters, which can be differentiated and updated.
More specifically, for the collection of the modules in the $i$-th layer, we utilize the row vector $\bar{\bm{\gamma}}^{i}\in[0,1]^{1\times N}$ of $\bm{\bar{\Gamma}}$ as learnable hyperparameters in continuous space. The forward pass of the relaxed LoRA-based fine-tuning formula can then be defined as:
$h_{j}^{i}=\bm{W}_{j}^{i}\bm{x}+\bar{\gamma}_{j}^{i}\Delta\bm{W}_{j}^{i}\bm{x}=\bm{W}_{j}^{i}\bm{x}+\bar{\gamma}_{j}^{i}\bm{B}_{j}^{i}\bm{A}_{j}^{i}\bm{x},$ (10)
where $h_{j}^{i}$ denotes the $j$-th item of the hidden nodes in the $i$-th layer, and each $\bar{\gamma}_{j}^{i}$ in $\bar{\bm{\gamma}}^{i}$ satisfies
$\bar{\gamma}_{j}^{i}=\frac{\exp(\gamma_{j}^{i})}{\sum_{j\in[N]}\exp(\gamma_{j}^{i})},$ (11)
$\bm{W}_{j}^{i}\in\mathbb{R}^{d\times k}$ is the pre-trained weight matrix of module $m_{j}^{i}$. $\bm{B}_{j}^{i}$ and $\bm{A}_{j}^{i}$ are $j$-th low-rank decomposition matrices in the $i$-th layer.
Intuitively, $\bm{\bar{\Gamma}}$ can represent the necessities of fine-tuning this module using $\mathbf{A}$ and $\mathbf{B}$. When this weight tends to $1$, it indicates that fine-tuning this module is necessary and might lead to better performance, and vice versa. By utilizing the continuous relaxation on the differential matrix, we are able to generate each $\bar{\bm{\gamma}}^{i}$ via a few rounds of training (e.g., five rounds, determined by the datasets). We denote $\mathcal{L}_{train}$ and $\mathcal{L}_{valid}$ as the training and validation loss, both of which are functions of $\Delta\bm{W}$ and $\bar{\bm{\Gamma}}$ such that
$\mathcal{L}_{\cdot}:=\mathcal{L}_{\cdot}(\Delta\bm{W},\bar{\bm{\Gamma}})$ (12)
Our goal is to find the best hyperparameters $\bar{\bm{\Gamma}}$ to minimize the validation loss, which is equivalent to bi-level optimization problems as follows:
$\min_{\bar{\bm{\Gamma}}}$ $\mathcal{L}_{valid}(\Delta\bm{W}^{\ast},\bar{\bm{\Gamma}})$ (13)
$s.t.$ $\Delta\bm{W}^{\ast}=\mathop{\arg\min}_{\Delta\bm{W}}\mathcal{L}_{train}(\Delta\bm{W},\bar{\bm{\Gamma}})$
To solve the optimizations in Eq. 13, we use the gradient descent algorithm to update the parameters $\bm{A}$, $\bm{B}$ and $\bar{\bm{\Gamma}}$. Specifically, we randomly extract part of the data in the training set and divide it into two parts, i.e., the training data and validation data, which are used to update $\Delta\bm{W}$ and $\bar{\bm{\Gamma}}$ alternately. We present the detailed training algorithms in Algorithm 1, lines 1 to 8. Note that during this step, the remaining network parameters are kept frozen as pre-trained.
4.2 Discretization and fine-tuning
Having the relaxed matrix $\bar{\bm{\Gamma}}$ in the previous stage, we perform the discretization to obtain the binary DAM $\bm{\Gamma}$. An intuitive representation of this procedure is shown in the right part of Fig. 1. More concretely, for each item in $\bar{\bm{\Gamma}}$, we update the top $K:=\lfloor\rho\cdot N\rfloor$ largest entries with 1 and set the remaining values as 0, such that
$\gamma^{i}_{j}=\begin{cases}1&\mbox{if $\bar{\gamma}^{i}_{j}\geq\delta^{i}$}\\ 0&\mbox{otherwise}\end{cases}$ (14)
where $\delta^{i}$ is the value of the $K$-th largest entry in $\bar{\bm{\gamma}}^{i}$ and $\rho$ is the selecting ratio. In our design, $\rho$ is a hyperparameter which is set as 0.5 for fair comparison with the baselines. The discretization is described in Algorithm 1, lines 10 and 11. The final step is to fine-tune the model equipped with the binarized DAM and low-rank decomposition matrices. During the fine-tuning, only the modules with weight one (e.g., the selected module) will be fine-tuned, while the others are kept unchanged in the downstream tasks.
4.3 Optimization and weight sharing
In the previous section, we obtain the DAM $\bm{\Gamma}$ using continuous relaxation and discretization. This approach can significantly improve the model performance after fine-tuning when the discrepancy of the weights in $\bm{\bar{\Gamma}}$ is distinct. However, when the entries in the relaxed $\bm{\bar{\Gamma}}$ display a uniform distribution, the method in the previous descriptions can be further optimized. More concretely, as shown in Fig. 2, we take the continuous DAM on CoLA and MRPC [31] as examples. We visualize matrix $\bar{\bm{\Gamma}}$ on these two datasets with DeBERTaV3-base [9] as the backbone, respectively, in which each column is a layer with six trainable modules. It can be seen that the weight distribution on CoLA is relatively distinct, i.e., in the majority of the layers, $W_{O}$ and $W_{Q}$ obtain significantly higher weights so that they are more suitable for the following fine-tuning. In contrast, the entries of the $\bm{\bar{\Gamma}}$ on MRPC share a similar and uniform distribution. For instance, all the candidate modules obtain weights around 0.15 from layers 1 to 8. Under this circumstance, if we select the top-$K$ largest entries, the modules that correspond to the “0” entries after discretization also demonstrate the same level of importance as those selected by ones. Consequently, directly fixing the modules with zero entries in $\mathbf{\Gamma}$ might lead to performance degradation.
Figure 2: The module weights of the DeBERTaV3-base model in the QNLI and RTE datasets. In the figure, $W_{Q}$, $W_{K}$, $W_{V}$, $W_{I}$, $W_{O}$ and $W_{D}$ correspond to the $\mathsf{query\_proj}$, $\mathsf{key\_proj}$, $\mathsf{value\_proj}$, $\mathsf{intermediate.dense}$, $\mathsf{output.dence}$, and $\mathsf{attention.output.dense}$ modules in the pre-trained model respectively.
To address this discretization discrepancy issue, we adopt the weight-sharing strategy to further optimize our method. Specifically, for the modules corresponding to the zero entries after discretization, instead of just freezing them without fine-tuning, we fine-tune them with the same weights as the modules in other layers. In other words, those modules with zero DAM entries share the same model weights in the fine-tuning procedure. For instance, all the modules $\bm{W}_{i}$ in layers 1 to 12 that are not selected will share the same weights and also participate in fine-tuning, where $\bm{W}_{i}\in\{\bm{W}_{Q},\bm{W}_{K},\bm{W}_{V},\bm{W}_{I},\bm{W}_{O},\bm{W}_{D}\}$. This optimization can enhance the performance of DiffoRA without introducing a large amount of extra fine-tuning parameters. The model fine-tuning with weight-sharing is shown in Algorithm 1, lines 12 to 15.
5 Experiments
5.1 Configurations
Hardware. DiffoRA is fully implemented in Python programming language. We evaluate our method on a Desktop Core i7-12700F CPU and GeForce RTX 3090.
Datasets and pre-trained models. We utilize two types of benchmarks: i) General Natural Language Understanding (GLUE) [30], including MNLI, SST-2, CoLA, QQP, QNLI, RTE, MRPC, and STS-B; and ii) Question Answering, including SQuADv1.1 [26] and SQuADv2.0 [27]. We use the DeBERTaV3-base [9] as the backbone model in the main text, and the results on GLUE using RoBERTa-base [20] are presented in Appendix D.
Counterpart comparisons. We use the following methods as baselines. (i) Full FT uses all parameters for fine-tuning; (ii) BitFit [35] is a sparse-fine-tuning method for pre-trained models that updates only a small subset of the bias terms; (iii) Houlsby adapter [11] adds a few trainable modules inserted between layers of a pre-trained model, allowing for task-specific tuning without altering the entire model; (iv) Pfeiffer adapter [24] combines multiple task-specific adapters by linearly blending their outputs; (v) LoRA [12] reduces the number of trainable parameters by applying low-rank matrix decomposition to weight updates in pre-trained models; (vi) AdaLoRA [36] adapts LoRA by dynamically adjusting the rank of low-rank updates during training, optimizing parameter efficiency while maintaining model performance across various tasks. We select the baseline methods with comparable parameter amounts and open-sourced codes for fair comparisons.
Implementation details. We set the module retention ratio $\rho$ to 50%, the LoRA rank to 4, and the $\alpha$ of LoRA to 16. All the results are the average values under three random seeds. See Appendix C for more detailed settings.
5.2 Natural language understanding
We evaluate the performance of the fine-tuned model on the GLUE benchmark [31] using various approaches. We use DeBERTaV3-base as the pre-trained model, which contains 183 million parameters. During this task, we set the rank of the low-rank decomposition matrices to 4 and fine-tuned 50% of the modules in each layer of the pre-trained model. We adopt all eight datasets in GLUE and the concrete fine-tuning model architectures are determined by the specific tasks. We summarize and present the results in Tab. 1. Overall, our DiffoRA achieves the highest model accuracy among the baselines on all the datasets. More concretely, our method achieves 0.81% higher fine-tuning accuracy than the state-of-the-art method AdaLoRA on the CoLA dataset. DiffoRA exhibits consistently better results than the existing methods under the same level of parameter amounts. Moreover, our method also outperforms the baseline methods with a larger parameter size, i.e., up to 3.84% higher accuracy than PAdapter with twice the parameters involved. Compared to the full fine-tuning strategy, DiffoRA obtains a better performance, which also demonstrates the effectiveness of our approach in the GLUE benchmark.
Method #Params MNLI SST-2 CoLA QQP QNLI RTE MRPC STS-B All
m/mm Acc Mcc Acc/F1 Acc Acc Acc Corr Avg.
Full FT 184M 89.90/90.12 95.63 69.19 92.40/89.80 94.03 83.75 89.46 91.60 88.09
BitFit 0.1M 89.37/89.91 94.84 66.96 88.41/84.95 92.24 78.70 87.75 91.35 86.02
HAdapter 0.61M 90.12/90.23 95.30 67.87 91.65/88.95 93.76 85.56 89.22 91.30 87.93
PAdapter 0.60M 90.15/90.28 95.53 69.48 91.62/88.86 93.98 84.12 89.22 91.52 88.04
HAdapter 0.31M 90.10/90.02 95.41 67.65 91.54/88.81 93.52 83.39 89.25 91.31 87.60
PAdapter 0.30M 89.89/90.06 94.72 69.06 91.40/88.62 93.87 84.48 89.71 91.38 87.90
LoRAr=2 0.33M 90.30/90.38 94.95 68.71 91.61/88.91 94.03 85.56 89.71 91.68 88.15
AdaLoRA 0.32M 90.09/90.41 95.80 70.04 91.78/89.16 94.49 87.36 90.44 91.63 88.81
DiffoRA 0.35M 90.49/90.49 96.09 70.85 91.79/89.12 94.52 87.96 90.79 91.75 89.11
Table 1: The results of the fine-tuned DeBERTaV3-base model on the GLUE dataset are presented, with the best results highlighted in bold and the second-best results underlined. Our DiffoRA achieved the best results on average.
5.3 Question answering
For the question-answering task, we use DeBERTaV3-base as the pre-trained model and adopt two datasets (i.e., SQuADv1.1 and SQuADv2.0) under different amounts of parameters to fine-tune the model. In order to keep the number of parameters close to the baseline, we choose the rank of the low-rank decomposition matrices from $\{1,2,5,10\}$ in the SQuADv1.1 dataset and $\{2,4,8,15\}$ in the SQuADv2.0 dataset. We use Exact Match (EM) and F1 as evaluation indicators and summarize the results in Tab. 2. Similar to the GLUE benchmark, our DiffoRA also demonstrates consistently better results than the baseline on both SQuAD datasets. Specifically, on SQuADv1.1, DiffoRA obtains 0.4% to 0.5% higher EM and 0.1% to 0.2% higher F1 than the best baseline AdaLoRA. Compared to the full fine-tuning method results, our DiffoRA also achieves higher accuracy even with 0.08% parameters fine-tuned. Furthermore, on the SQuADv2.0 dataset, our method is around 0.2% higher than the best baseline on EM and F1. DiffoRA is consistently better than the fully fine-tuning strategy (i.e., 85.4% EM and 88.4% F1) and original LoRA, which demonstrates the effectiveness of our scheme.
In all, the experimental results in this section demonstrate that DiffoRA works consistently better than the baseline methods on all benchmarks and datasets, which conforms to our theoretical analysis.
Method SQuADv1.1 SQuADv2.0
Full FT 86.0/92.7 85.4/88.4
#Params 0.08% 0.16% 0.32% 0.65% 0.08% 0.16% 0.32% 0.65%
HAdapter 84.4/91.5 85.3/92.1 86.1/92.7 86.7/92.9 83.4/86.6 84.3/87.3 84.9/87.9 85.4/88.3
PAdapter 84.4/91.7 85.9/92.5 86.2/92.8 86.6/93.0 84.2/87.2 84.5/87.6 84.9/87.8 84.5/87.5
LoRA 86.4/92.8 86.6/92.9 86.7/93.1 86.7/93.1 83.0/86.3 83.6/86.7 84.5/87.4 85.0/88.0
AdaLoRA 87.2/93.4 87.5/93.6 87.5/93.7 87.6/93.7 83.0/86.3 84.6/87.5 84.1/87.3 84.2/87.3
DiffoRA 87.6/93.5 88.1/93.8 88.1/93.8 88.1/93.9 84.2/87.2 84.8/87.8 85.1/88.0 85.5/88.4
Table 2: The results of the fine-tuned DeBERTaV3-base model on the SQuAD dataset. We report EM/F1. The best results are highlighted in bold and the second-best results are underlined.
6 Analysis
DiffoRA v.s. Random select strategy. We first analyze the effectiveness of DiffoRA by comparing its performance with that of the random selection method. Specifically, we randomly select three modules in each layer as the fine-tuning modules, and the remaining modules are processed with the weight-sharing strategy. The results are shown in Tab. 3. We choose ranks from $\{1,2,4\}$ for fine-tuning in the STS-B and SQuADv1.1 datasets, respectively. It can be demonstrated that compared to random sampling, DiffoRA achieves significant performance improvements and effectively identifies important modules, e.g., DiffoRA is 0.2% to 0.6% higher than the random select strategy.
Method Rank #Params STS-B SQuADv1.1
Random 1 0.09 91.21 87.23
DiffoRA 1 0.08 91.65 87.37
Random 2 0.19 91.38 87.55
DiffoRA 2 0.16 91.52 87.74
Random 4 0.37 91.11 87.82
DiffoRA 4 0.34 91.75 88.04
Table 3: Random Selection v.s. DiffoRA on two datasets.
Sample rate. We select sample rates from {0.2, 0.4, 0.5, 0.7, 0.9} corresponding to the most important Top-$\{1,2,3,4,5\}$ modules, respectively. To explore the relationship between the performance and sample rate, we conduct experiments on three datasets: STS-B, RTE, and MRPC. The results are summarized in Tab. 4. The results show that when the sampling rate is around 0.5, the performance of the fine-tuned model achieves state-of-the-art.
Sample Rate K STS-B RTE SQuADv1.1
0.2 1 90.87 85.07 87.09
0.4 2 91.66 85.65 87.43
0.5 3 91.75 87.96 88.12
0.7 4 91.41 87.04 88.19
0.9 5 91.39 86.34 88.18
Table 4: DiffoRA across three datasets at different sample rates.
Weight sharing. We further investigate the effects of the weight-sharing strategy on our DiffoRA. Specifically, we consider two scenarios: (i) weight sharing, and (ii) weight sharing combined with a selection matrix. Experiments are conducted on the STS-B and SQuADv1.1 datasets, and the results are summarized in Tab. 5. The table shows that the combination of module selection and weight sharing can lead to better results. For instance, our DiffoRA with the weight sharing achieves 0.11% to 2.9% higher accuracy than the method without this strategy.
Weight Share STS-B RTE MRPC
91.75 87.96 90.79
91.66 85.06 89.58
Table 5: Comparison of DiffoRA w/o weight sharing.
7 Conclusion
We propose a new PEFT method called DiffoRA, which enables efficient and adaptive LLM fine-tuning based on LoRA. Instead of adjusting every interior rank, we argue that adopting LoRA module-wisely is sufficient. To achieve this, we construct a DAM to select the modules that are most suitable and essential to fine-tune. We theoretically analyze how the DAM impacts the convergence rate and generalization capability. Furthermore, we adopt continuous relaxation and discretization to establish DAM. To alleviate the issue of discretization discrepancy, we utilize the weight-sharing strategy for optimization. The experimental results demonstrate that our DiffoRA works consistently better than the baselines across all benchmarks.
Supplementary Material
Appendix A Detailed Proofs
A.1 Proof of Theorem 2
We first introduce the Wely inequality as follows.
Lemma 1 (Weyl inequality [10]).
Let $\bm{A}$, $\bm{B}\in\mathbb{R}^{n\times n}$ be Hermitian matrices, and let the eigenvalues of $\bm{A}$, $\bm{B}$, and $\bm{A}+\bm{B}$ be $\{\lambda_{i}(\bm{A})\}_{i=1}^{n}$, $\{\lambda_{i}(\bm{B})\}_{i=1}^{n}$ and $\{\lambda_{i}(\bm{A}+\bm{B})\}_{i=1}^{n}$, respectively. The eigenvalues of each matrix are arranged in ascending order. Then we have
$\lambda_{i}(\bm{A}+\bm{B})\leq\lambda_{i+j}(\bm{A})+\lambda_{n-j}(\bm{B}),\quad j=\{0,1,\ldots,n-i\}$ (15)
for each $i\in[n]$, with equality for some pair $i,j$ if and only if there is a nonzero vector $\bm{x}$ such that $\bm{A}\bm{x}=\lambda_{i+j}(\bm{A})\bm{x}$, $\bm{B}\bm{x}=\lambda_{n-j}(\bm{B})\bm{x}$, and $(\bm{A}+\bm{B})\bm{x}=\lambda_{i}(\bm{A}+\bm{B})\bm{x}$. Also,
$\lambda_{i-j+1}(\bm{A})+\lambda_{j}(\bm{B})\leq\lambda_{i}(\bm{A}+\bm{B}),\quad j=\{1,\ldots,i\}$ (16)
for each $i\in[n]$, with equality for some pair $i,j$ if and only if there is a nonzero vector $\bm{x}$ such that $\bm{A}\bm{x}=\lambda_{i-j+1}(\bm{A})\bm{x}$, $\bm{B}\bm{x}=\lambda_{j}(\bm{B})\bm{x}$, and $(\bm{A}+\bm{B})\bm{x}=\lambda_{i}(\bm{A}+\bm{B})\bm{x}$. If $\bm{A}$ and $\bm{B}$ have no common eigenvector, then inequality (15) and (16) are strict inequality.
We first present and proof the following lemma.
Lemma 2.
If $\bm{x}_{i}\nparallel\bm{x}_{j},\forall i\neq j$, we have $\bm{H}^{\infty}_{w_{0}}\succ 0$.
Proof.
By the Lemma 3.4 of [4], there exists $\bm{w}\sim\mathcal{N}(\bm{0},\bm{I})$, such that when $m$ is sufficiently large, $\|\bm{w}-\bm{w}_{0}\|$ is sufficiently small. Then according to the proof of Theorem 3.1 of [4], we get $\lambda_{\min}(\bm{H}_{w_{0}}^{\infty})>0$.
Then we provide the proof of Theorem 2.
Theorem 2. Suppose $f$ is an NN with a single hidden layer and ReLU activation function. Assume $\bm{X}\in\mathbb{R}^{d\times n}$, $\mathbf{w}(0)\sim N(\mathbf{0},\mathbf{I})$, hidden nodes $m=\Omega\left(\frac{n^{6}d^{2}}{(\lambda_{0}^{\Gamma})^{4}\delta^{3}}\right)$, and $\bm{I}^{\bm{\Gamma w}}-\bm{I}^{\bm{w}}\succeq 0$, then the following formula holds with probability at least $1-\delta$ over the initialization
$\|f(\mathbf{W}(t),\mathbf{a},\mathbf{X};\bm{\Gamma},\bm{W}_{0})-\bm{y}\|_{2}^{2}$ (17)
$\leq$ $\exp(-\lambda_{0}^{\Gamma}t)\|f(\mathbf{W}(0),\mathbf{a},\mathbf{X};\bm{\Gamma},\bm{W}_{0})-\bm{y}\|_{2}^{2}$
where $\lambda_{0}^{\Gamma}\geq\lambda_{0}$.
Proof.
We denote $\bm{I}^{-}:=\bm{I}^{\bm{\Gamma w}}-\bm{I}^{\bm{w}}\succeq 0$. Then we have the following inequalities:
$\lambda_{\min}(\bm{H}^{\infty}_{\Gamma,w_{0}})$ $=\lambda_{\min}(\bm{H}^{\infty}_{w_{0}}+\bm{H}^{\infty}_{\Gamma,w_{0}}-\bm{H}_{w_{0}}^{\infty})$ (18)
$\geq\lambda_{\min}(\bm{H}_{w_{0}}^{\infty})+\lambda_{\min}(\bm{H}^{\infty}_{\Gamma,w_{0}}-\bm{H}_{w_{0}}^{\infty})$
From the definitions of $\bm{H}^{\infty}_{\Gamma,w_{0}}$ and $\bm{H}^{\infty}_{w_{0}}$ we have
$\bm{H}^{\infty}_{\Gamma,w_{0}}-\bm{H}_{w_{0}}^{\infty}=\bm{X}^{T}\bm{X}\odot\bm{I}^{-}$ (19)
Since $\bm{X}^{T}\bm{X}$ and $\bm{I}^{-}$ are both positive definite/semi-positive definite matrices, their Hadamard product is also a positive definite/semi-positive definite matrix [22], i.e., $\bm{H}^{\infty}_{\Gamma,w_{0}}-\bm{H}^{\infty}_{w_{0}}\succeq 0$. Therefore, we have
$\lambda_{\min}(\bm{H}_{\Gamma,w_{0}}^{\infty})\geq\lambda_{\min}(\bm{H}_{w_{0}}^{\infty})$ (20)
Finally, according to Theorem 1, we have
$\|f(\mathbf{W}(t),\mathbf{a},\mathbf{X};\bm{\Gamma},\bm{W}_{0})-\bm{y}\|_{2}^{2}$ (21)
$\leq$ $\exp(-\lambda_{0}^{\Gamma}t)\|f(\mathbf{W}(0),\mathbf{a},\mathbf{X};\bm{\Gamma},\bm{W}_{0})-\bm{y}\|_{2}^{2}$
where $\lambda_{0}^{\Gamma}\geq\lambda_{0}$. ∎
Datasets learning rate batch size epochs share r r K $\alpha$ LoRA drop warm-up early stop
MNLI 3e-4 196 7 2 4 3 16 0.25 4000 -
RTE 7e-4 32 50 2 4 3 16 0.2 900 -
QNLI 6e-4 64 5 1 6 3 16 0.35 3000 -
MRPC 1e-3 16 14 2 4 3 16 0 900 -
QQP 8e-4 64 5 2 4 3 16 0 2000 -
SST-2 1e-4 32 5 2 4 3 16 0 3000 -
CoLA 3e-4 16 6 2 4 3 16 0.15 900 5
STS-B 7e-4 16 7 2 4 3 16 0 900 -
Table 6: Training settings for GLUE benchmarks.
Settings SQuADv1.1 SQuADv2.0
#Params 0.08% 0.16% 0.32% 0.65% 0.08% 0.16% 0.32% 0.65%
r 1 2 5 10 2 4 8 15
train epochs 3 3
learning rate 5e-4 7e-4
warm-up 2000 4000
share r 1 2
K 3 3
$\alpha$ 16 16
LoRA drop 16 16
Table 7: Training settings for SQuAD benchmarks.
A.2 Proof of Theorem 3
Theorem 3. For an over-parameterized neural network with the loss on the testing set as $\mathcal{L}(\bm{W},\bm{a};\bm{\Gamma},\bm{W}_{0})$. Let $\bm{y}=(y_{1},...,y_{N})^{T}$, and $\eta=\kappa C_{1}\sqrt{\bm{y}^{T}(\bm{H}_{\Gamma,w_{0}}^{\infty})^{-1}\bm{y}}/(m\sqrt{N})$ for some small enough absolute constant $\kappa$, where $\eta$ denotes the step of SGD. Under the assumption of Theorem 2, for any $\delta\in(0,e^{-1}]$, there exists $m^{\ast}(\delta,N,\lambda_{0}^{\Gamma})$, such that if $m\geq m^{*}$, then with probability at least $1-\delta$, we have
$\mathbb{E}[\mathcal{L}(\bm{W},\bm{a};\bm{\Gamma},\bm{W}_{0})]\leq$ $\mathcal{O}(C^{\prime}\sqrt{\frac{\bm{y}^{T}\bm{y}}{\lambda_{0}^{\Gamma}N}})+\mathcal{O}(\sqrt{\frac{\log(1/\delta)}{N}})$ (22)
where $\lambda_{0}^{\Gamma}\geq\lambda_{0}$, $C,C^{\prime}$, and $\delta$ are constants.
Proof.
According to the courant minimax principle [5], D.2 in [39], we get
$\bm{y}^{T}(\mathbf{H}_{\Gamma,w_{0}}^{\infty})^{-1}\bm{y}\leq\frac{\bm{y}^{T}\bm{y}}{\lambda_{\min}(\mathbf{H}_{\Gamma,w_{0}}^{\infty})}.$
Thus, we have
$\mathbb{E}[\mathcal{L}(\bm{W},\bm{a};\bm{\Gamma},\bm{W}_{0})]$
$\leq O\left(c\cdot\sqrt{\frac{\bm{y}^{T}(\mathbf{H}_{\Gamma,w_{0}}^{\infty})^{-1}\bm{y}}{N}}\right)+O\left(\sqrt{\frac{\log(1/\delta)}{N}}\right)$
$\leq O\left(c\cdot\sqrt{\frac{\bm{y}^{T}\bm{y}}{N\lambda^{\Gamma}_{0}}}\right)+O\left(\sqrt{\frac{\log(1/\delta)}{N}}\right)$
Method #Params MNLI SST-2 CoLA QQP QNLI RTE MRPC STS-B All
Acc Acc Mcc Acc Acc Acc Acc Corr Avg.
Full FT 124.65M 87.68 94.73 60.26 90.75 92.58 78.63 88.33 90.31 85.41
BitFit 0.1M 85.50 94.38 61.16 88.08 90.99 79.57 89.07 90.55 84.91
HAdapter 1.20M 86.53 93.73 62.62 90.83 92.82 80.43 89.90 90.16 85.88
PAdapter 1.19M 86.75 93.83 63.87 90.53 92.61 80.51 89.51 90.65 86.03
LoRA 1.33M 87.11 93.71 63.54 90.44 92.76 80.65 89.90 90.91 86.13
AdaLoRA 1.27M 87.89 95.11 63.23 90.48 92.84 81.23 89.02 91.22 86.38
FLoRA 1.33M 87.43 94.27 63.31 90.38 92.75 81.59 90.44 90.82 86.37
DiffoRA 1.32M 87.73 95.16 64.95 91.04 92.84 81.98 89.44 91.35 86.81
Table 8: The results of the fine-tuned RoBerta-base model on the GLUE dataset are presented, with the best results highlighted in bold and the second-best results underlined.
Appendix B Dataset Details
GLUE [31] The GLUE Benchmark is a comprehensive collection of natural language understanding tasks, designed to evaluate the performance of models across various NLP applications. It includes:
  • MNLI: Multinomial Natural Language Inference (inference task), including 393k training data and 20k test data.
  • SST-2: Stanford Sentiment Treebank (sentiment analysis task), including 67k training data and 1.8k test data.
  • MRPC: Microsoft Research Paraphrase Corpus (paraphrase detection task), including 3.7k training data and 1.7k test data.
  • CoLA: Corpus of Linguistic Acceptability (linguistic acceptability task), including 8.5k training data and 1k test data.
  • QNLI: Question Natural Language Inference (inference task), including 108k training data and 5.7k test data.
  • QQP: Quora Question Pairs (question-answering task), including 364k training data and 391k test data.
  • RTE: Recognizing Textual Entailment (inference task), including 7k training data and 1.4k test data.
  • STS-B: Semantic Textual Similarity Benchmark (textual similarity), including 7k training data and 1.4k test data.
SQuAD SQuADv1.1 [26] is a dataset consisting of approximately 100k question-answer pairs based on a collection of Wikipedia articles. The task is to extract exact spans of text from the articles as answers to the given questions. SQuADv2.0 [27] builds on v1.1 by adding unanswerable questions. It includes about 150k question-answer pairs from over 500 articles, requiring models to both extract answers and identify when no answer is available.
Appendix C Training Details
We summarize the training settings of GLUE and SQuAD in Tab. 6 and Tab. 7, respectively. In order to compare with the baseline method at the same parameter level, we use different $r$ for different datasets. We use “share r” to represent the rank of the shared modules.
Appendix D Results on RoBERTa-base
Finally, to further prove the effectiveness of our proposed DiffoRA, we adopt RoBERTa-base as the backbone model which contains 125 million parameters. We fine-tune the pre-trained model on eight tasks of the GLUE benchmark. We set the rank of the low-rank decomposition matrices to $4\sim 12$ and fine-tuned $20\%\sim 70\%$ of the modules in each layer of the pre-trained model. The experimental results are summarized in Tab. 8. It can be shown that our DiffoRA achieves the best results in almost all datasets. For instance, our approach outperforms the state-of-the-art method PAdapter on the CoLA dataset by 0.98%. Although DiffoRA achieves the second-best result on a few tasks, we achieve the highest average accuracy among all the baselines, i.e., 86.81% on average. The experimental results demonstrate the effectiveness of our method.