### v2

# DiffoRA: Enabling Parameter-Efficient Fine-Tuning via Differential Module Selection

Abstract
The Parameter-Efficient Fine-Tuning (PEFT) methods have been extensively researched for large language models in downstream tasks. Among all the existing approaches, the Low-Rank Adaptation (LoRA) has gained popularity for its streamlined design by incorporating low-rank matrices into existing pre-trained models. Though effective, LoRA, as well as its adaptive optimizations, either allocate the same matrix to all the modules or adjust the interior rank of the components based on importance scoring indicators. In this paper, we argue that not all the modules in LLMs are suitable and necessary to be fine-tuned. Enlightened by this insight, we propose a new PEFT scheme called DiffoRA, which enables adaptive adoption of the low-rank decomposition matrices. At the core of DiffoRA lies a Differential Adaptation Matrix (DAM) to determine which module is the most suitable and essential for fine-tuning. We theoretically explain how the designed matrix impacts the convergence rate and generalization capability of a pre-trained model. We then construct the DAM via continuous relaxation and discretization with weight-sharing optimizations. We fully implement DiffoRA and design comprehensive experiments to evaluate its performance. The experimental results demonstrate that DiffoRA delivers state-of-the-art results across multiple benchmarks.
1 Introduction
Large language models (LLMs) have gained significant attention across various domains of natural language processing, demonstrating remarkable capabilities in tasks such as text generation [1, 2], machine translation [3], sentiment analysis [4], and question-answering [5]. Due to the large model size, fine-tuning an LLM to downstream tasks will invoke a large number of parameters, e.g., up to 175 billion parameters to fully fine-tune GPT-3. As a result, Parameter-Efficient Fine-Tuning (PEFT) has become increasingly paramount.
The existing PEFT methods can be categorized into two classes. Some approaches slightly modify the model structure by adding small trainable modules and keeping the rest of the model unchanged [6, 7]. Another line of work aims to efficiently capture the incremental parameter updates without modifying the model structure [8]. Among all the existing work, the Low-Rank Adaptation [9] is widely acknowledged due to its effectiveness and satisfying performance. Unlike the previous work, LoRA incurs some low-rank decomposition matrices to parameterize the incremental updates and realizes comparable results with 70% less overhead than the full fine-tuning. This innovative approach paves the way for more effective utilization of LLMs in practice.
Although LoRA is effective in practice, it treats all the modules in the network equally by adding the same decomposition matrices to all the trainable modules in the network. To alleviate this issue, some work has been proposed as adaptive LoRA, which managed to adjust the interior ranks of decomposition matrices for each module based on parameter importance estimation [10, 11, 12, 13, 14]. For example, Zhang et al. [11] utilizes the singular value of the low-rank matrices to adjust the rank, while SalientLoRA [15] uses a salience measurement to adaptively optimize the ranks of LoRA.
Though effective, we point out that fine-tuning all modules, like the previous work, is not necessary and might inversely affect the model performance. As justification, we conduct a toy experiment by comparing the accuracy of DeBERTaV3-Base on CoLA [16] using LoRA when selectively fine-tuning: S1) all modules; S2) only the QKV projections; S3) only dense modules; S4) QKV projections in the first half of the model and dense modules in the remaining half; and S5) three modules per layer that are randomly selected. We repeat the selection for 10 times to find the optimal configuration. It can be seen from the results in Figure 1 that
  • Different modules and combination strategies contribute differently to the fine-tuning results. An appropriate module combination can significantly enhance the model’s performance. For instance, by randomly selecting the modules in each layer, S5 achieves over 4% higher accuracy than other settings with lower parameter amounts.
  • Not all of the modules are suitable and necessary to be fine-tuned. For instance, S1 obtains a lower accuracy than S4 and S5, which also leads to overfitting. Some modules in S1 contribute negatively to the performance, thus, it is unnecessary to adjust their interior ranks.
The above observations demonstrate the importance of selecting the most essential modules during LoRA fine-tuning, which is often overlooked by the existing methods. This presents a compelling research question: Can we construct a theoretically grounded PEFT method to enable adaptive adoption of the low-rank decomposition matrices, so that we can only fine-tune the modules that are most necessary and essential?
Figure 1: The loss (left) and final accuracy (right) obtained on the CoLA dataset and DeBERTaV3-base using LoRA under different settings.
Our contributions. In this paper, we affirmatively answer the question with DiffoRA, a novel method to improve PEFT performance based on LoRA. Instead of adjusting the interior rank of every decomposition matrix, DiffoRA only adopts the low-rank matrices for those modules that are suitable and necessary to be fine-tuned. To this end, at the core of our approach lies a Differentiable Adaptation Matrix (DAM) that performs a module selection. DAM is designed to be differentiable with respect to the incremental updates of the low-rank matrices, allowing it to effectively capture fine-tuning priorities for various network backbones. We provide a theoretical analysis of how DAM influences convergence rate and generalization during fine-tuning. Then, we approximate the NP-hard optimization by constructing DAM via continuous relaxation and discretization. To mitigate discretization-induced discrepancies and enhance robustness, we further introduce a weight-sharing strategy. We implement DiffoRA and conduct extensive experiments on three standard benchmarks for both Natural Language Understanding and Question Answering tasks. The results demonstrate that DiffoRA consistently outperforms existing methods. For example, it achieves a 1.51% accuracy gain over the state-of-the-art on the CoLA dataset. The main contributions of this work are summarized as follows:
  • We propose DiffoRA, a novel and theoretically-grounded PEFT method for LLMs. Our approach is built atop a newly designed differentiable matrix (i.e., DAM), which enables module-wise adoption of the low-rank decomposition matrices.
  • We theoretically explain how the DAM impacts the performance of a LoRA-based fine-tuning model in terms of the convergence rate and generalization capability.
  • We fully implement DiffoRA and evaluate its performance on several widely adopted benchmarks that contain multiple tasks. The experimental results demonstrate that DiffoRA works consistently better than all the baselines.
2 Related work
2.1 Parameter-efficient fine-tuning.
LLMs have captured considerable public interest due to their success in multiple regions. The PEFT is essential for LLMs due to the huge number of parameters. Some previous work has been proposed to fine-tune the LLMs using specifically designed modules that are added to the LLMs. The fine-tuning of LLMs is thus converted to the adjustments of these small modules. For instance, multiple methods [17, 6, 7, 18] insert dataset-dependent small modules or vectors between the layers to decrease the parameter amounts. Another line of work models the incremental updates of the fine-tuning procedure to make it parameter-efficient. Guo et al. [8] propose to use a task-specific difference vector to extend and fine-tune the base model. There are also some methods that fine-tune parts of the parameters [19], e.g., the bias of FFN [20] or the last quarter [21].
2.2 Low-rank adaptation and optimizations
To further reduce the computational and storage cost, Hu et al. [9] proposed LoRA, in which they designed a low-rank decomposition matrix to model the incremental updates during fine-tuning. A plethora of work has been proposed to optimize LoRA [22, 23, 24, 25, 26, 27, 28, 29]. One of the limitations of LoRA is that it treats all the modules equally, which omits the variations of the modules in LLMs. To address this issue, a few work has been proposed to realize an adaptive LoRA. These methods first evaluate the importance of the modules and then adjust the interior rank of the decomposition matrices accordingly, e.g., AdaLoRA [11] and SalientLoRA [15]. There is also a few work that utilize the training procedures to determine the module importance [30, 10, 12], yet they still focus on modifying the interior ranks of the decomposition matrices.
3 Theoretical analysis
3.1 Preliminaries
Notations. We define $[n]=\{1,2,\ldots,n\}$. Vectors and matrices are represented by bold lowercase and bold uppercase fonts, respectively. For instance, $\bm{x}$ is a vector with entry $x_{i}$, and $\bm{M}$ is a matrix with entry $[\bm{M}]_{ij}$. The minimum eigenvalue of $\bm{M}$ is denoted as $\lambda_{\min}(\bm{M})$. $\|\cdot\|_{2}$ is used to represent the $l_{2}$ norm of a vector. $N(\bm{0},\bm{I})$ and $U\{S\}$ represent the standard Gaussian distribution and uniform distribution over a set $S$, respectively. We denote by $\bm{X}=\{(\bm{x}_{i},y_{i})|\bm{x}_{i}\in\mathbb{R}^{d\times 1},y_{i}\in\mathbb{R},i\in[n]\}$ the training set, where $\bm{x}_{i}$ and $y_{i}$ represent the $i$-th data and label. $\mathbb{I}\{\cdot\}$ represents the indicator function that demonstrates the event occurrence, such that for event $\mathcal{A}$, $\mathbb{I}\{\mathcal{A}\}=1$ if and only if $\mathcal{A}$ happened, otherwise it equals 0. $P(\mathcal{A})$ represents the probability of $\mathcal{A}$ occurred event.
For input $\bm{x}\in\mathbb{R}^{d\times 1}$, weight vector $\bm{w}\in\mathbb{R}^{d\times 1}$ in the weight matrix $\bm{W}\in\mathbb{R}^{d\times m}$, and output weight $\bm{a}\in\mathbb{R}^{m\times 1}$, we denote $f(\bm{W},\bm{a},\bm{x})$ as a neural network with a single hidden layer such that
$f(\bm{W},\bm{a},\bm{x})=\frac{1}{\sqrt{m}}\sum\nolimits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}^{T}\bm{x})$ (1)
where $\sigma$ is the activation function. In this paper, we primarily consider the ReLU function, which is one of the most widely adopted activation functions in the literature, i.e., $\sigma(z)=z\mathbb{I}\{z>0\}$. Given a training set $\bm{X}$, the optimization goal is to minimize the empirical risk loss function
$\mathcal{L}(\bm{W},\bm{a})=\sum\nolimits_{i=1}^{n}\frac{1}{2}(f(\bm{W},\bm{a},\bm{x}_{i})-y_{i})^{2}$ (2)
Recall LoRA. Hu et al. [9] utilize two matrices $\bm{A}\in\mathbb{R}^{r\times k},\bm{B}\in\mathbb{R}^{d\times r}$ to substitute the parameters’ increments, where $d,k\ll r$, $\bm{A}$ is usually initialized by following Gaussian distribution and $\bm{B}$ is initialized with zeros. For $\bm{h}=\bm{W}^{0}\bm{x}$, the forward pass in LoRA is
$\bm{h}=\bm{W}^{0}\bm{x}+\Delta\bm{x}=\bm{W}^{0}\bm{x}+\bm{B}\cdot\bm{A}\bm{x}$ (3)
where $\bm{W}^{0}$ denotes the weight matrix of the pre-trained network, consisting of $r$ row vectors $\bm{w}_{0,r}$. In the following, we denote $\Delta\bm{W}$ as $\bm{B}\cdot\bm{A}$. LoRA adopts this modification equally to all the modules in the model. Therefore, fine-tuning the pre-trained network using LoRA can be formulated as
$f(\bm{W},\bm{a},\bm{x};\bm{W}_{0})=\frac{1}{\sqrt{m}}\sum\nolimits_{r=1}^{m}a_{r}\sigma((\bm{w}_{0,r}+\bm{w}_{r})^{T}\bm{x})$ (4)
where $\bm{w}_{r}$ is the $r$-th row of the trainable parameters in the low-rank matrices.
3.2 Main theorems
Technical intuitions. In this work, we aim to construct a matrix called DAM to determine the necessity of each module for fine-tuning in LLM. We denote DAM as $\bm{\Gamma}$. Intuitively, $\mathbf{\bm{\Gamma}}$ is applied to the trainable parameters of the LoRA module-wise, such that the module corresponding to $\gamma_{i,j}=1$ is determined to be fine-tuned, and vice versa. Notably, the module-wise fine-tuning can be viewed as a special case of an element-wise formulation, where entire submatrices are either enabled or disabled uniformly. For analytical clarity, we focus on the more general element-wise setting, which allows finer-grained control and provides deeper insights into how the structure of $\bm{\Gamma}$ influences models’ performance. In what follows, we assume access to a suitable algorithm for constructing $\bm{\Gamma}$ and aim to elucidate how such selective adaptation contributes to improved model performance. To facilitate this analysis, we extend the definitions introduced in Equation 1 as follows:
$f(\bm{W},\bm{a},\bm{x};\bm{\Gamma},\bm{W}_{0})=\frac{1}{\sqrt{m}}\sum_{r=1}^{m}a_{r}\sigma((\bm{w}_{0,r}+\bm{\gamma}_{r}\circ\bm{w}_{r})^{T}\bm{x})$ (5)
where $\bm{\gamma}_{r}$ is the $r$-th row of $\bm{\Gamma}$ and $\circ$ represents the Hadamard product. Furthermore, we follow the definitions in [31] and define $\mathbf{H}^{\infty}_{\bm{\Gamma},\bm{w}_{0}}$ and $\mathbf{H}^{\infty}_{\bm{w}_{0}}$ based on $\bm{\Gamma}$ such that
Definition 1 (Gram Matrix with $\bm{\Gamma}$).
For a neural network with a single hidden layer and ReLU activation, the Gram matrix $\bm{H}^{\infty}_{\bm{\Gamma},\bm{w}_{0}}\in\mathbb{R}^{n\times n}$ induced on a training set $\bm{X}:=\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$ is defined entry-wise as
$[\bm{H}_{\bm{\Gamma},\bm{w}_{0}}^{\infty}]_{ij}=$ $\mathbb{E}_{\bm{w}\sim N(\bm{0},\bm{I}),r\in[m]}[\bm{x}_{i}^{T}\bm{x}_{j}\cdot\mathbb{I}\{(\bm{w}_{0,r}+\bm{\gamma}_{r}\circ\bm{w})^{T}\bm{x}_{i}\geq 0,(\bm{w}_{0,r}+\bm{\gamma}_{r}\circ\bm{w})^{T}\bm{x}_{j}\geq 0\}]$ (6)
$=$ $\bm{x}_{i}^{T}\bm{x}_{j}[\bm{I}^{\bm{\Gamma}\bm{w}}]_{ij}$
We also construct $\mathbf{H}_{\bm{w}_{0}}^{\infty}$ with entry $[\mathbf{H}_{\bm{w}_{0}}^{\infty}]_{ij}$ such that
$[\bm{H}^{\infty}_{\bm{w}_{0}}]_{ij}=$ $\mathbb{E}_{\bm{w}\sim N(\bm{0},\bm{I}),r\in[m]}[\bm{x}_{i}^{T}\bm{x}_{j}\cdot\mathbb{I}\{(\bm{w}_{0,r}+\bm{w})^{T}\bm{x}_{i}\geq 0,(\bm{w}_{0,r}+\bm{w})^{T}\bm{x}_{j}\geq 0\}]$ (7)
$=$ $\bm{x}_{i}^{T}\bm{x}_{j}[\bm{I}^{\bm{w}}]_{ij}$
where $\bm{I}^{\bm{\bm{\Gamma}w}}$ and $\bm{I}^{\bm{w}}$ are the expectations of the indicator matrices. We denote $\lambda_{0}:=\lambda_{\min}(\mathbf{H}_{\bm{w}_{0}}^{\infty})$, and $\lambda_{0}^{\bm{\Gamma}}:=\lambda_{\min}(\mathbf{H}_{\bm{\Gamma},\bm{w}_{0}}^{\infty})$.
Theoretical results. The theoretical analysis is based on the intuitive observation that fine-tuning a pre-trained model can be viewed as training a well-initialized model. We analyze LoRA fine-tuning from the perspective of over-parameterized neural networks, focusing on two fundamental aspects: the convergence rate and the generalization capability. Over-parameterized networks have been widely adopted and demonstrated strong empirical performance in hierarchical feature extraction, primarily due to their large capacity and training tractability [32, 33]. In this section, we leverage this architecture to establish our theoretical understanding of the matrix $\mathbf{\bm{\Gamma}}$. Building on prior work [31], we present the following theorem that characterizes the convergence behavior of a single-hidden-layer neural network in the over-parameterized setting.
We first analyze the relationship of the minimum eigenvalue between the adaptive matrix (i.e., $\lambda^{\bm{\Gamma}}_{0}$) and the original matrix (i.e., $\lambda_{0}$). We summarize our results in the following theorem:
Theorem 1.
Suppose $f$ is an NN with a single hidden layer and ReLU activation function. Assume $\bm{X}\in\mathbb{R}^{d\times n}$, $\bm{w}(0)\sim N(\bm{0},\bm{I})$, hidden nodes $m=\Omega\left(\frac{n^{6}d^{2}}{(\lambda_{0}^{\bm{\Gamma}})^{4}\delta^{3}}\right)$, and $\bm{I}^{\bm{\bm{\Gamma}w}}-\bm{I}^{\bm{w}}\succeq 0$, then the following formula holds with probability at least $1-\delta$ over the initialization
$\|f(\bm{W}(t),\bm{a},\bm{X};\bm{\Gamma},\bm{W}_{0})-\bm{y}\|_{2}^{2}$ $\leq\exp(-\lambda_{0}^{\bm{\Gamma}}t)\|f(\bm{W}(0),\bm{a},\bm{X};\bm{\Gamma},\bm{W}_{0})-\bm{y}\|_{2}^{2}$ (8)
$\leq\exp(-\lambda_{0}t)\|f(\bm{W}(0),\bm{a},\bm{X};\bm{\Gamma},\bm{W}_{0})-\bm{y}\|_{2}^{2}$
where $\lambda_{0}^{\bm{\Gamma}}\geq\lambda_{0}$.
Proof sketch. The key to the proof lies in identifying relationship between $\lambda_{0}$ and $\lambda_{0}^{\bm{\Gamma}}$ by Weyl inequalities [34]. We provide the full proof in Section A.1.
Theorem 1 demonstrates that the minimum eigenvalue of the Gram matrix of the network with $\mathbf{\bm{\Gamma}}$ is larger than the $\lambda_{0}$ (without the selective matrix), thus leading to a higher convergence rate. Other than the convergence rate, we also analyze the relationship between $\lambda_{0}^{\bm{\Gamma}}$ and the generalization capability of the model. We present the results in the following theorem:
Theorem 2.
For an over-parameterized neural network with the loss on the testing set as $\mathcal{L}(\bm{W},\bm{a};\bm{\Gamma},\bm{W}_{0})$. Let $\bm{y}=(y_{1},...,y_{N})^{T}$, and $\eta=\kappa C\sqrt{\bm{y}^{T}(\bm{H}_{\bm{\Gamma},\bm{w}_{0}}^{\infty})^{-1}\bm{y}}/(m\sqrt{N})$ for some small enough absolute constant $\kappa$, where $\eta$ denotes the step of SGD. Under the assumption of Theorem 1, for any $\delta\in(0,e^{-1}]$, there exists $m^{\ast}(\delta,N,\lambda_{0}^{\bm{\Gamma}})$, such that if $m\geq m^{*}$, then with probability at least $1-\delta$, we have
$\mathbb{E}[\mathcal{L}(\bm{W},\bm{a};\bm{\Gamma},\bm{W}_{0})]$ $\leq\mathcal{O}(C^{\prime}\sqrt{\frac{\bm{y}^{T}\bm{y}}{\lambda_{0}^{\bm{\Gamma}}N}})+\mathcal{O}(\sqrt{\frac{\log(1/\delta)}{N}})\leq\mathcal{O}(C^{\prime}\sqrt{\frac{\bm{y}^{T}\bm{y}}{\lambda_{0}N}})+\mathcal{O}(\sqrt{\frac{\log(1/\delta)}{N}})$ (9)
where $\mathcal{L}(\bm{W},\bm{a};\bm{\Gamma},\bm{W}_{0})$ is the loss corresponding to Equation 5, $C,C^{\prime}$, and $\delta$ are constants.
Proof sketch. The proof of the above theorem derives from Corollary 3.10 of [35] and Section D.2 of [36]. We present the detailed proof in Section A.2.
Similarly to Theorem 1, the above theorem indicates that the adoption of $\bm{\Gamma}$ can enhance the generalization capability of the model, as shown in the last inequality of Equation 9.
Remark 1.
In this section, we theoretically demonstrate that incorporating the DAM matrix $\bm{\Gamma}$ into a LoRA-based network can improve both the convergence rate and generalization performance, ultimately leading to enhanced model effectiveness. Theorem 1 and 2 provide formal justifications for the benefits of $\bm{\Gamma}$. The remaining challenge lies in how to efficiently construct such a matrix. We address this in the following section, where we introduce practical solutions and detailed algorithms.
4 Design of DiffoRA
We now present the concrete method to construct the module-wise selective matrix $\mathbf{\bm{\Gamma}}$. The proposed approach is called DiffoRA, which is built atop the DAM (i.e., Differentiable Adaptation Matrix). To capture the information of the incremental updates modeled by the low-rank decomposition matrices, DiffoRA views the elements $\gamma$ in $\bm{\Gamma}$ as trainable parameters. More concretely, as shown in Figure 2, DiffoRA approximates the NP-hard problem by invoking the following two stages: (i) Relaxation and optimization, which aims to map $\bm{\Gamma}$ to a continuous space (i.e., $\bar{\bm{\Gamma}}$) so that it is differentiable and can be further optimized; and (ii) Discretization and fine-tuning, which binarizes $\bar{\bm{\Gamma}}$ to determine the most essential module for fine-tuning. We will first elaborate on each stage, respectively, and then introduce the optimization method called weight-sharing.
Figure 2: The framework of DiffoRA contains two stages. In stage one (left part), the initialized adaptive matrix is continuously relaxed, i.e., C.R. in the figure. The DAM is then differentiable and can be updated. In stage two (right part), the obtained DAM is discretized into binary. All the modules corresponding to entry one will be fine-tuned using decomposition matrices.
4.1 Continuous relaxation
The first stage of DiffoRA is illustrated in the left part of Figure 2. In contrast with the existing work, we point out that it is unnecessary to allocate ranks for every module. Instead, we can construct a selective matrix called DAM that determines which module needs to be fine-tuned and which does not. The final output $\bm{\Gamma}$ is a binary matrix, in which the “ones” (resp. “zeros”) indicate that the corresponding entries will (resp. will not) be fine-tuned. However, directly generating a binary matrix is non-trivial, i.e., it is an NP-hard problem. To address this challenge, DiffoRA relaxes the binary constraint on the DAM matrix from $\bm{\Gamma}\in\{0,1\}^{L\times N}$ to a continuous form $\bar{\bm{\Gamma}}\in[0,1]^{L\times N}$, where $L$ and $N$ denote the number of model layers and the number of trainable modules per layer, respectively. All row vectors in $\bar{\bm{\Gamma}}$ are treated as differentiable hyperparameters, allowing them to be jointly optimized with the network via gradient-based methods.
More specifically, for the collection of the modules in the $i$-th layer, we utilize the row vector $\bar{\bm{\gamma}}_{i}\in[0,1]^{1\times N}$ of $\bm{\bar{\bm{\Gamma}}}$ as learnable hyperparameters in continuous space. The forward pass of the relaxed LoRA-based fine-tuning formula can then be defined as:
$h_{i,j}=\bm{W}_{i,j}\bm{x}+\bar{\gamma}_{i,j}\Delta\bm{W}_{i,j}\bm{x}=\bm{W}_{i,j}\bm{x}+\bar{\gamma}_{i,j}\bm{B}_{i,j}\bm{A}_{i,j}\bm{x},$ (10)
where $h_{i,j}$ denotes the $j$-th item of the hidden nodes in the $i$-th layer, and each $\bar{\gamma}_{i,j}$ in $\bar{\bm{\gamma}}_{i}$ satisfies $\sum_{j}\bar{\gamma}_{i,j}=1$, and $\bar{\gamma}_{i,j}\geq 0$. $\bm{W}_{i,j}\in\mathbb{R}^{d\times k}$ is the pre-trained weight matrix of module $m_{i,j}$. $\bm{B}_{i,j}$ and $\bm{A}_{i,j}$ are the $j$-th low-rank decomposition matrices in the $i$-th layer.
Intuitively, $\bm{\bar{\bm{\Gamma}}}$ can represent the necessities of fine-tuning this module using $\mathbf{A}$ and $\mathbf{B}$. When this weight tends to $1$, it indicates that fine-tuning this module is necessary and might lead to better performance, and vice versa. By utilizing the continuous relaxation on the differential matrix, we are able to generate each $\bar{\bm{\gamma}}_{i}$ via a few rounds of training (e.g., five rounds, determined by the datasets). We denote $\mathcal{L}_{train}$ and $\mathcal{L}_{valid}$ as the training and validation loss. Then, our goal is to find the best hyperparameters $\bar{\bm{\Gamma}}$ to minimize the validation loss, which is equivalent to bi-level optimization problems as follows:
$\min_{\bar{\bm{\Gamma}}}$ $\mathcal{L}_{valid}(\Delta\bm{W}^{\ast},\bar{\bm{\Gamma}})$ (11)
$s.t.$ $\Delta\bm{W}^{\ast}=\mathop{\arg\min}_{\Delta\bm{W}}\mathcal{L}_{train}(\Delta\bm{W},\bar{\bm{\Gamma}})$
To solve Equation 11, we use the gradient descent algorithm to update the parameters $\bm{A}$, $\bm{B}$ and $\bar{\bm{\Gamma}}$. Specifically, we randomly extract part of the data in the training set and divide it into two parts, i.e., the training data and validation data, which are used to update $\Delta\bm{W}$ and $\bar{\bm{\Gamma}}$ alternately. We present the detailed training algorithms in Algorithm 1 in Appendix E, lines 1 to 8.
4.2 Discretization and fine-tuning
Having the relaxed matrix $\bar{\bm{\Gamma}}$ in the previous stage, we perform the discretization to obtain the binary DAM $\bm{\Gamma}$. An intuitive representation of this procedure is shown in the right part of Figure 2. More concretely, for each item in $\bar{\bm{\Gamma}}$, we update the top $K:=\lfloor\rho\cdot N\rfloor$ largest entries with 1 and set the remaining values as 0, such that
$\gamma_{i,j}=\begin{cases}1&\mbox{if $\bar{\gamma}_{i,j}\geq\delta_{i}$}\\ 0&\mbox{otherwise}\end{cases}$ (12)
where $\delta_{i}$ is the value of the $K$-th largest entry in $\bar{\bm{\gamma}}_{i}$ and $\rho$ is the selection ratio. The discretization is described in Algorithm 1, lines 10 and 11. The final step is to fine-tune the model equipped with the binarized DAM and low-rank decomposition matrices. During the fine-tuning, only the modules with weight one (e.g., the selected module) will be fine-tuned, while the others are kept unchanged.
4.3 Optimization with weight sharing
Figure 3: The module weights of the DeBERTaV3-base model in the CoLA and SST-2 datasets.
In the previous section, we constructed the DAM $\bm{\Gamma}$ via continuous relaxation and discretization. While this method significantly improves fine-tuning performance when the relaxed weights $\bar{\bm{\Gamma}}$ are highly discriminative, its effectiveness diminishes under uniformly distributed weights. As shown in Figure 3, we visualize $\bar{\bm{\Gamma}}$ on CoLA and SST-2 [16] using DeBERTaV3-base [37] as the backbone, where each column represents a layer with six candidate modules. On CoLA, weight distributions are relatively distinct—e.g., $W_{Q}$ dominates in most layers, facilitating effective module selection. In contrast, SST-2 exhibits nearly uniform weights, with all modules receiving values around 0.15 across layers 8–12. In such cases, top-$K$ selection may discard modules that are equally important, leading to potential performance degradation if zero-valued modules in $\bm{\Gamma}$ are entirely frozen.
To address this issue, we introduce a weight-sharing optimization strategy. For insufficiently discriminative DAM, we do not exclude modules with zero entries from the fine-tuning process, but allow them to share weights with corresponding modules in different layers. Specifically, all unselected modules $\bm{W}_{i}\in\{\bm{W}_{Q},\bm{W}_{K},\bm{W}_{V},\bm{W}_{I},\bm{W}_{O},\bm{W}_{D}\}$ for all layers share the same set of weights and are jointly updated during fine-tuning. This strategy is particularly beneficial for handling DAMs with ambiguous distinctions, enabling modules with low discriminability to still participate in optimization. As a result, DiffoRA’s effectiveness is enhanced without introducing additional parameter overhead. The corresponding implementation details are provided in Algorithm 1, lines 12–15.
5 Experiments
5.1 Configurations
Datasets and pre-trained models. We utilize two types of benchmarks: i) General Natural Language Understanding (GLUE) [38], including MNLI, SST-2, CoLA, QQP, QNLI, RTE, MRPC, and STS-B; and ii) Question Answering, including SQuADv1.1 [39] and SQuADv2.0 [40]. In the main text, we adopt DeBERTaV3-base [37], LLaMA-3.2-1B [41], and LLaMA-7B [42] as backbone models. Additionally, results obtained on the GLUE benchmark using RoBERTa-base [43] are provided in Appendix F for comparison.
Counterpart comparisons. We choose representative LoRA and adaptive LoRA methods from the literature as baselines, ensuring comparable parameter budgets and availability of open-source implementations. The details can be found in Appendix B. Our code is available at the https://anonymous.4open.science/r/DiffoRA-2128.
Table 1: The results of the fine-tuned DeBERTaV3-base model on the GLUE dataset are presented, with the best results highlighted in bold and the second-best results underlined.
Method #Params MNLI SST-2 CoLA QQP QNLI RTE MRPC STS-B All
m/mm Acc Mcc Acc/F1 Acc Acc Acc Corr Avg.
Full FT 184M 89.90/90.12 95.63 69.19 92.40/89.80 94.03 83.75 89.46 91.60 88.09
BitFit 0.1M 89.37/89.91 94.84 66.96 88.41/84.95 92.24 78.70 87.75 91.35 86.02
HAdapter 0.61M 90.12/90.23 95.30 67.87 91.65/88.95 93.76 85.56 89.22 91.30 87.93
PAdapter 0.60M 90.15/90.28 95.53 69.48 91.62/88.86 93.98 84.12 89.22 91.52 88.04
HAdapter 0.31M 90.10/90.02 95.41 67.65 91.54/88.81 93.52 83.39 89.25 91.31 87.60
PAdapter 0.30M 89.89/90.06 94.72 69.06 91.40/88.62 93.87 84.48 89.71 91.38 87.90
LoRAr=2 0.33M 90.30/90.38 94.95 68.71 91.61/88.91 94.03 85.56 89.71 91.68 88.15
AdaLoRA 0.32M 90.66/90.70 95.80 70.04 91.78/89.16 94.49 87.36 90.44 91.63 88.86
DiffoRA 0.35M 90.55/90.55 96.09 71.55 92.07/89.46 94.59 88.57 91.42 91.86 89.66
±$\pm 0.01$/±$\pm 0.01$$\pm 0.00$$\pm 1.33$$\pm 0.02$/±$\pm 0.03$$\pm 0.08$$\pm 0.52$$\pm 0.88$$\pm 0.04$ $\uparrow 0.80$
Table 2: The results of the fine-tuned DeBERTaV3-base model on the SQuAD dataset. We report EM/F1. The best results are highlighted in bold and the second-best results are underlined.
Method SQuADv1.1 SQuADv2.0
Full FT 86.0/92.7 85.4/88.4
#Params 0.08% 0.16% 0.32% 0.65% 0.08% 0.16% 0.32% 0.65%
HAdapter 84.4/91.5 85.3/92.1 86.1/92.7 86.7/92.9 83.4/86.6 84.3/87.3 84.9/87.9 85.4/88.3
PAdapter 84.4/91.7 85.9/92.5 86.2/92.8 86.6/93.0 84.2/87.2 84.5/87.6 84.9/87.8 84.5/87.5
LoRA 86.4/92.8 86.6/92.9 86.7/93.1 86.7/93.1 83.0/86.3 83.6/86.7 84.5/87.4 85.0/88.0
AdaLoRA 87.2/93.4 87.5/93.6 87.5/93.7 87.6/93.7 83.0/86.3 84.6/87.5 84.1/87.3 84.2/87.3
DiffoRA 87.6/93.5 88.1/93.8 88.1/93.8 88.1/93.9 84.2/87.2 84.8/87.8 85.1/88.0 85.5/88.4
5.2 Natural language understanding
We evaluate the performance of the fine-tuned model on the GLUE benchmark [16]. We use DeBERTaV3-base as the pre-trained model, which contains 183 million parameters. During this task, we set the rank of the low-rank decomposition matrices to 4 and fine-tuned 50% of the modules in each layer. We adopt all eight datasets in GLUE, and the concrete fine-tuning model architectures are determined by the specific tasks. We summarize and present the results in Table 1. Overall, our DiffoRA model achieves the highest accuracy among all baselines at the same parameter count. More specifically, on the CoLA dataset, our method outperforms AdaLoRA by 1.51% in fine-tuning accuracy, and 0.8% on average across all tasks. In addition, we observe that DiffoRA consistently outperforms LoRA and the adaptive LoRA method, suggesting that some modules do not need to be fine-tuned, which is consistent with our observations.
5.3 Question answering
For the question-answering task, we fine-tune DeBERTaV3-base on SQuADv1.1 and SQuADv2.0 with varying parameter budgets. To ensure comparability with baselines, we set the LoRA rank to ${1,2,5,10}$ for SQuADv1.1 and ${2,4,8,15}$ for SQuADv2.0. Evaluation is conducted using Exact Match (EM) and F1, with results summarized in Table 2. DiffoRA consistently outperforms all baselines across both datasets. On SQuADv1.1, it achieves 0.4–0.5% higher EM and 0.1–0.2% higher F1 than AdaLoRA. Remarkably, it also surpasses full fine-tuning while updating only 0.08% of the parameters. Similar gains are observed on SQuADv2.0, where DiffoRA outperforms all baselines and even full fine-tuning, further validating its effectiveness.
5.4 Incorporating DiffoRA with existing methods
To show the effectivenss of DiffoRA, we further incorporate it with LoRA+ [27], which is an optimized LoRA method. We evaluate the performance on LLaMA-3.2-1B and LLaMA-7B across three NLU benchmarks: MNLI, SST-2, and QQP.. As shown in Table 3, both DiffoRALoRA and DiffoRALoRA+ consistently outperform their counterparts. For example, on MNLI, DiffoRA improves the accuracy of LLaMA-3.2-1B by 2.25% and 2.61% over LoRA and LoRA+, respectively, while using fewer tunable parameters. These results confirm the superiority of DiffoRA over both standard and optimized LoRA variants across model scales.
In summary, the experimental results presented in this section demonstrate that DiffoRA consistently outperforms all baseline methods across various benchmarks and datasets. This highlights the importance of selecting the right modules for fine-tuning and aligns well with our theoretical insights.
Table 3: The results of the fine-tuned LLaMA model on the MNLI, SST-2 and QQP datasets are presented, with the best results highlighted in bold and the second-best results underlined.
Model Method #Params (%) MNLI (Acc) SST-2 (Acc) QQP (Acc/F1) All (Avg.)
LLaMA-3.2-1B LoRA 0.45 83.94 95.30 90.39/86.89 89.13
LoRA+ 0.45 84.80 92.43 88.21/84.19 87.40
[…]
LLaMA-7B LoRA 0.30 90.74 97.02 90.75/87.22 91.43
LoRA+ 0.30 89.71 95.99 90.88/87.67 91.06
[…]
6 Further analysis
DiffoRA vs. Random selection strategy. We analyze the effectiveness of DiffoRA by comparing its performance with that of the random selection method. Specifically, we randomly select three modules in each layer as the fine-tuning modules, and the remaining modules are processed with the weight-sharing strategy. The results are shown in Table 5. We choose ranks from $\{1,2,4\}$ for fine-tuning in the STS-B and SQuADv1.1 datasets, respectively. It can be demonstrated that compared to random sampling, DiffoRA achieves significant performance improvements and effectively identifies important modules, e.g., DiffoRA is 0.2% to 0.6% higher than the random select strategy.
Sample rate. We study the effects of different sample rates on DiffoRA. To explore the relationship between the performance and sample rate, we conduct experiments on three datasets: STS-B, RTE, and SQuADv1.1. We select sample rates from {0.2, 0.4, 0.5, 0.7, 0.9} corresponding to the most important Top-$\{1,2,3,4,5\}$ modules, respectively. The results are summarized in Table 5. The results show that when the sampling rate is around 0.5, the performance of the fine-tuned model achieves state-of-the-art.
Table 4: Random Selection vs. DiffoRA.
Method Rank #Params STS-B SQuADv1.1
Random 1 0.07 91.23 87.23
DiffoRA 1 0.07 91.38 87.37
Random 2 0.14 91.49 87.55
DiffoRA 2 0.13 91.81 87.74
Random 4 0.27 90.99 87.82
DiffoRA 4 0.34 91.86 88.04
Table 5: DiffoRA across three datasets at different sample rates.
Sample Rate K STS-B RTE SQuADv1.1
0.2 1 91.34 87.73 87.09
0.4 2 91.86 84.84 87.43
0.5 3 91.41 88.57 88.12
0.7 4 91.18 88.45 88.19
0.9 5 91.58 86.64 88.18
7 Conclusion and limitations
We introduced DiffoRA, a novel and theoretically-grounded PEFT method that enables module-wise LoRA adoption. Instead of adjusting the interior ranks, DiffoRA employs a selective matrix DAM to identify essential modules, backed by theoretical analysis on convergence and generalization. The DAM is constructed through continuous relaxation and discretization, with a weight-sharing strategy to mitigate discretization discrepancies. Experiments across diverse benchmarks show that DiffoRA consistently outperforms prior methods. Currently, the sample rate is determined by empirical observations, which differs across models and datasets. This might lead to suboptimal results. In future work, we aim to investigate principled strategies for sample rate selection to improve generalization and adaptability of our method.
Appendix A Detailed Proofs
A.1 Proof of Theorem 1
We first introduce the main theorem in [31] and Wely inequality as follows.
Theorem 3 (Theorem 3.3 in [31]).
If gram matrix $\bm{H}^{\infty}\succ 0$, $\|\bm{x}_{i}\|_{2}=1$, $|y_{i}|<C$ for some constant $C$ and $i\in[n]$, hidden nodes $m=\Omega\left(\frac{n^{6}}{\lambda_{\min}(\bm{H}^{\infty})^{4}\delta^{3}}\right)$, and i.i.d. initialize $\bm{w}_{r}\sim N(\bm{0},\bm{I})$, $a_{r}\sim U\{[-1,1]\}$ for $r\in[m]$, then with probability at least $1-\delta$ over the initialization, the following inequality holds:
$\|f(\bm{W}(t),\bm{a},\bm{X})-\bm{y}\|_{2}^{2}\leq\exp{(-\lambda_{\min}(\bm{H}^{\infty})t)}\|f(\bm{W}(0),\bm{a},\bm{X})-\bm{y}\|_{2}^{2}$ (13)
where
$\bm{H}^{\infty}:=\mathbb{E}_{\bm{w}\sim N(\bm{0},\bm{I})}[\bm{x}_{i}^{T}\bm{x}_{j}\mathbb{I}\{\bm{w}^{T}\bm{x}_{i}\geq 0,\bm{w}^{T}\bm{x}_{j}\geq 0\}].$
The inequality in the above theorem demonstrates that the minimum eigenvalue of the Gram matrix positively affects the training convergence rate of the network.
Lemma 1 (Weyl inequality [34]).
Let $\bm{A}$, $\bm{B}\in\mathbb{R}^{n\times n}$ be Hermitian matrices, and let the eigenvalues of $\bm{A}$, $\bm{B}$, and $\bm{A}+\bm{B}$ be $\{\lambda_{i}(\bm{A})\}_{i=1}^{n}$, $\{\lambda_{i}(\bm{B})\}_{i=1}^{n}$ and $\{\lambda_{i}(\bm{A}+\bm{B})\}_{i=1}^{n}$, respectively. The eigenvalues of each matrix are arranged in ascending order. Then we have
$\lambda_{i}(\bm{A}+\bm{B})\leq\lambda_{i+j}(\bm{A})+\lambda_{n-j}(\bm{B}),\quad j=\{0,1,\ldots,n-i\}$ (14)
for each $i\in[n]$, with equality for some pair $i,j$ if and only if there is a nonzero vector $\bm{x}$ such that $\bm{A}\bm{x}=\lambda_{i+j}(\bm{A})\bm{x}$, $\bm{B}\bm{x}=\lambda_{n-j}(\bm{B})\bm{x}$, and $(\bm{A}+\bm{B})\bm{x}=\lambda_{i}(\bm{A}+\bm{B})\bm{x}$. Also,
$\lambda_{i-j+1}(\bm{A})+\lambda_{j}(\bm{B})\leq\lambda_{i}(\bm{A}+\bm{B}),\quad j=\{1,\ldots,i\}$ (15)
for each $i\in[n]$, with equality for some pair $i,j$ if and only if there is a nonzero vector $\bm{x}$ such that $\bm{A}\bm{x}=\lambda_{i-j+1}(\bm{A})\bm{x}$, $\bm{B}\bm{x}=\lambda_{j}(\bm{B})\bm{x}$, and $(\bm{A}+\bm{B})\bm{x}=\lambda_{i}(\bm{A}+\bm{B})\bm{x}$. If $\bm{A}$ and $\bm{B}$ have no common eigenvector, then inequality (14) and (15) are strict inequality.
We first present and proof the following lemma.
Lemma 2.
If $\bm{x}_{i}\nparallel\bm{x}_{j},\forall i\neq j$, we have $\bm{H}^{\infty}_{\bm{w}_{0}}\succ 0$.
Proof.
By the Lemma 3.4 of [31], there exists $\bm{w}\sim\mathcal{N}(\bm{0},\bm{I})$, such that when $m$ is sufficiently large, $\|\bm{w}-\bm{w}_{0}\|$ is sufficiently small. Then according to the proof of Theorem 3.1 of [31], we get $\lambda_{\min}(\bm{H}_{\bm{w}_{0}}^{\infty})>0$.
Then we provide the proof of Theorem 2.
Theorem 2. Suppose $f$ is an NN with a single hidden layer and ReLU activation function. Assume $\bm{X}\in\mathbb{R}^{d\times n}$, $\mathbf{w}(0)\sim N(\mathbf{0},\mathbf{I})$, hidden nodes $m=\Omega\left(\frac{n^{6}d^{2}}{(\lambda_{0}^{\bm{\Gamma}})^{4}\delta^{3}}\right)$, and $\bm{I}^{\bm{\bm{\Gamma}w}}-\bm{I}^{\bm{w}}\succeq 0$, then the following formula holds with probability at least $1-\delta$ over the initialization
$\|f(\mathbf{W}(t),\mathbf{a},\mathbf{X};\bm{\Gamma},\bm{W}_{0})-\bm{y}\|_{2}^{2}$ (16)
$\leq$ $\exp(-\lambda_{0}^{\bm{\Gamma}}t)\|f(\mathbf{W}(0),\mathbf{a},\mathbf{X};\bm{\Gamma},\bm{W}_{0})-\bm{y}\|_{2}^{2}$
where $\lambda_{0}^{\bm{\Gamma}}\geq\lambda_{0}$.
Proof.
We denote $\bm{I}^{-}:=\bm{I}^{\bm{\bm{\Gamma}w}}-\bm{I}^{\bm{w}}\succeq 0$. Then we have the following inequalities:
$\lambda_{\min}(\bm{H}^{\infty}_{\bm{\Gamma},\bm{w}_{0}})$ $=\lambda_{\min}(\bm{H}^{\infty}_{\bm{w}_{0}}+\bm{H}^{\infty}_{\bm{\Gamma},\bm{w}_{0}}-\bm{H}_{\bm{w}_{0}}^{\infty})$ (17)
$\geq\lambda_{\min}(\bm{H}_{\bm{w}_{0}}^{\infty})+\lambda_{\min}(\bm{H}^{\infty}_{\bm{\Gamma},\bm{w}_{0}}-\bm{H}_{\bm{w}_{0}}^{\infty})$
From the definitions of $\bm{H}^{\infty}_{\bm{\Gamma},\bm{w}_{0}}$ and $\bm{H}^{\infty}_{\bm{w}_{0}}$ we have
$\bm{H}^{\infty}_{\bm{\Gamma},\bm{w}_{0}}-\bm{H}_{\bm{w}_{0}}^{\infty}=\bm{X}^{T}\bm{X}\odot\bm{I}^{-}$ (18)
Since $\bm{X}^{T}\bm{X}$ and $\bm{I}^{-}$ are both positive definite/semi-positive definite matrices, their Hadamard product is also a positive definite/semi-positive definite matrix [44], i.e., $\bm{H}^{\infty}_{\bm{\Gamma},\bm{w}_{0}}-\bm{H}^{\infty}_{\bm{w}_{0}}\succeq 0$. Therefore, we have
$\lambda_{\min}(\bm{H}_{\bm{\Gamma},\bm{w}_{0}}^{\infty})\geq\lambda_{\min}(\bm{H}_{\bm{w}_{0}}^{\infty})$ (19)
Finally, according to Theorem 1, we have
$\|f(\mathbf{W}(t),\mathbf{a},\mathbf{X};\bm{\Gamma},\bm{W}_{0})-\bm{y}\|_{2}^{2}$ (20)
$\leq$ $\exp(-\lambda_{0}^{\bm{\Gamma}}t)\|f(\mathbf{W}(0),\mathbf{a},\mathbf{X};\bm{\Gamma},\bm{W}_{0})-\bm{y}\|_{2}^{2}$
where $\lambda_{0}^{\bm{\Gamma}}\geq\lambda_{0}$. ∎
Table 6: Training settings for GLUE benchmarks.
Datasets learning rate batch size epochs r K $\alpha$ LoRA drop warm-up
MNLI 3e-4 64 7 4 3 16 0.25 4000
RTE 5e-4 16 50 4 3 16 0.1 200
QNLI 6e-4 64 5 6 3 16 0.35 3000
MRPC 5e-4 32 30 4 3 16 0 200
QQP 4e-4 32 20 4 3 16 0 2000
SST-2 1e-4 32 5 4 3 16 0 3000
CoLA 2e-4 8 20 4 3 16 0 200
STS-B 5e-4 16 20 4 2 16 0.2 300
Table 7: Training settings for SQuAD benchmarks.
Settings SQuADv1.1 SQuADv2.0
#Params 0.08% 0.16% 0.32% 0.65% 0.08% 0.16% 0.32% 0.65%
r 1 2 5 10 2 4 8 15
train epochs 3 3
learning rate 5e-4 7e-4
warm-up 2000 4000
share r 1 2
K 3 3
$\alpha$ 16 16
LoRA drop 16 16
A.2 Proof of Theorem 2
Theorem 2. For an over-parameterized neural network with the loss on the testing set as $\mathcal{L}(\bm{W},\bm{a};\bm{\Gamma},\bm{W}_{0})$. Let $\bm{y}=(y_{1},...,y_{N})^{T}$, and $\eta=\kappa C_{1}\sqrt{\bm{y}^{T}(\bm{H}_{\bm{\Gamma},\bm{w}_{0}}^{\infty})^{-1}\bm{y}}/(m\sqrt{N})$ for some small enough absolute constant $\kappa$, where $\eta$ denotes the step of SGD. Under the assumption of Theorem 1, for any $\delta\in(0,e^{-1}]$, there exists $m^{\ast}(\delta,N,\lambda_{0}^{\bm{\Gamma}})$, such that if $m\geq m^{*}$, then with probability at least $1-\delta$, we have
$\mathbb{E}[\mathcal{L}(\bm{W},\bm{a};\bm{\Gamma},\bm{W}_{0})]\leq$ $\mathcal{O}(C^{\prime}\sqrt{\frac{\bm{y}^{T}\bm{y}}{\lambda_{0}^{\bm{\Gamma}}N}})+\mathcal{O}(\sqrt{\frac{\log(1/\delta)}{N}})$ (21)
where $\lambda_{0}^{\bm{\Gamma}}\geq\lambda_{0}$, $C,C^{\prime}$, and $\delta$ are constants.
Proof.
According to the courant minimax principle [45], D.2 in [36], we get
$\bm{y}^{T}(\mathbf{H}_{\bm{\Gamma},\bm{w}_{0}}^{\infty})^{-1}\bm{y}\leq\frac{\bm{y}^{T}\bm{y}}{\lambda_{\min}(\mathbf{H}_{\bm{\Gamma},\bm{w}_{0}}^{\infty})}.$
Thus, we have
$\mathbb{E}[\mathcal{L}(\bm{W},\bm{a};\bm{\Gamma},\bm{W}_{0})]$
$\leq O\left(c\cdot\sqrt{\frac{\bm{y}^{T}(\mathbf{H}_{\bm{\Gamma},\bm{w}_{0}}^{\infty})^{-1}\bm{y}}{N}}\right)+O\left(\sqrt{\frac{\log(1/\delta)}{N}}\right)$
$\leq O\left(c\cdot\sqrt{\frac{\bm{y}^{T}\bm{y}}{N\lambda^{\bm{\Gamma}}_{0}}}\right)+O\left(\sqrt{\frac{\log(1/\delta)}{N}}\right)$
Table 8: The results of the fine-tuned RoBerta-base model on the GLUE dataset are presented, with the best results highlighted in bold and the second-best results underlined.
Method #Params MNLI SST-2 CoLA QQP QNLI RTE MRPC STS-B All
Acc Acc Mcc Acc Acc Acc Acc Corr Avg.
Full FT 124.65M 87.68 94.73 60.26 90.75 92.58 78.63 88.33 90.31 85.41
BitFit 0.1M 85.50 94.38 61.16 88.08 90.99 79.57 89.07 90.55 84.91
HAdapter 1.20M 86.53 93.73 62.62 90.83 92.82 80.43 89.90 90.16 85.88
PAdapter 1.19M 86.75 93.83 63.87 90.53 92.61 80.51 89.51 90.65 86.03
LoRA 1.33M 87.11 93.71 63.54 90.44 92.76 80.65 89.90 90.91 86.13
AdaLoRA 1.27M 87.89 95.11 63.23 90.48 92.84 81.23 89.02 91.22 86.38
FLoRA 1.33M 87.43 94.27 63.31 90.38 92.75 81.59 90.44 90.82 86.37
DiffoRA 1.32M 87.73 95.16 64.95 91.04 92.84 81.98 89.44 91.35 86.81
Appendix B Baseline details
We use the following methods as baselines. (i) Full FT uses all parameters for fine-tuning; (ii) BitFit [20] is a sparse-fine-tuning method for pre-trained models that updates only a small subset of the bias terms; (iii) Houlsby adapter [46] adds a few trainable modules inserted between layers of a pre-trained model, allowing for task-specific tuning without altering the entire model; (iv) Pfeiffer adapter [47] combines multiple task-specific adapters by linearly blending their outputs; (v) LoRA [9] reduces the number of trainable parameters by applying low-rank matrix decomposition to weight updates in pre-trained models; (vi) AdaLoRA [11] adapts LoRA by dynamically adjusting the rank of low-rank updates during training, optimizing parameter efficiency while maintaining model performance across various tasks. (vii) LoRA+ [27] employs different learning rates to update the low-rank matrices.
Appendix C Dataset details
GLUE [16] The GLUE Benchmark is a comprehensive collection of natural language understanding tasks, designed to evaluate the performance of models across various NLP applications. It includes:
  • MNLI: Multinomial Natural Language Inference (inference task), including 393k training data and 20k test data.
  • SST-2: Stanford Sentiment Treebank (sentiment analysis task), including 67k training data and 1.8k test data.
  • MRPC: Microsoft Research Paraphrase Corpus (paraphrase detection task), including 3.7k training data and 1.7k test data.
  • CoLA: Corpus of Linguistic Acceptability (linguistic acceptability task), including 8.5k training data and 1k test data.
  • QNLI: Question Natural Language Inference (inference task), including 108k training data and 5.7k test data.
  • QQP: Quora Question Pairs (question-answering task), including 364k training data and 391k test data.
  • RTE: Recognizing Textual Entailment (inference task), including 7k training data and 1.4k test data.
  • STS-B: Semantic Textual Similarity Benchmark (textual similarity), including 7k training data and 1.4k test data.
SQuAD SQuADv1.1 [39] is a dataset consisting of approximately 100k question-answer pairs based on a collection of Wikipedia articles. The task is to extract exact spans of text from the articles as answers to the given questions. SQuADv2.0 [40] builds on v1.1 by adding unanswerable questions. It includes about 150k question-answer pairs from over 500 articles, requiring models to both extract answers and identify when no answer is available.
Appendix D Training details
DiffoRA is fully implemented in Python. We evaluate the method on a Desktop Core i7-12700F CPU and Tesla V100 GPU. We summarize the training settings of GLUE and SQuAD in Table 6 and Table 7, respectively. In order to compare with the baseline method at the same parameter level, we use different $r$ for different datasets.
Appendix E Algorithm pseudocode
We present the training algorithm in Algorithm 1.
Algorithm 1 DiffoRA
0:  Pre-trained model with $L$ layers $\mathcal{M}^{L}$; Candidate fine-tuning modules $M$, where $|M|=L\times N$, $N$ is the number of candidate modules in each layers; Training dataset and valid dataset $\bm{X}_{train}$ and $\bm{X}_{valid}$; Training epochs and valid epochs, $T$ and $V$; The learning rate $\eta$; Sample rate $\rho$; LoRA rank: $r_{l}$; Share rank: $r_{s}$.
1:  // Stage 1: Continuous Relaxation
2:  Create the hyperparameters $\bar{\bm{\Gamma}}\in\mathbb{R}^{L\times N}$.
3:  for $v=1;v<V;v++$ do
4:     Update hyperparameters $\bar{\bm{\Gamma}}$ as $\bar{\bm{\Gamma}}_{v}=\bar{\bm{\Gamma}}_{v}-\eta\nabla_{\bar{\bm{\Gamma}}}\mathcal{L}_{valid}(\bm{X}_{valid};\Delta\bm{W},\bar{\bm{\Gamma}}_{v-1})$
5:     for $t=1;t<T;t++$ do
6:        Update low-rank matrix $\Delta\bm{W}$ as follow:
$\Delta\bm{W}=\Delta\bm{W}-\eta\nabla_{\Delta\bm{W}}\mathcal{L}_{train}(\bm{X}_{train};\Delta\bm{W},\bar{\bm{\Gamma}}_{v})$
7:     end for
8:  end for
9:  // Stage 2: Discretization and Fine-Tuning
10:  Select the Top-K modules of each layer in $\mathcal{M}^{L}$ according to the $\bar{\bm{\Gamma}}$, where $K=\lfloor\rho\cdot N\rfloor$.
11:  Add a low-rank matrix $\Delta\bm{W}=\bm{BA}$ to the selected module and add the weight sharing matrix $\Delta\bm{W}_{s}=\bm{B}_{s}\bm{A}_{s}$ to the remaining modules, where $\bm{B}^{T},\bm{A}\in\mathbb{R}^{d\times r_{l}}$, and $\bm{B}^{T}_{s},\bm{A}_{s}\in\mathbb{R}^{d\times r_{s}}$.
12:  for $t=1;t<T;t++$ do
13:     Update low-rank matrix $\Delta\bm{W}$ as follow:
$\Delta\bm{W}=\Delta\bm{W}-\eta\nabla_{\Delta\bm{W}}\mathcal{L}_{train}(\bm{X}_{train};\Delta\bm{W})$
14:  end for
15:  return  Fine-tuned model $\mathcal{M}^{L}$.
Appendix F Results on RoBERTa-base
Finally, to further prove the effectiveness of our proposed DiffoRA, we adopt RoBERTa-base as the backbone model which contains 125 million parameters. We fine-tune the pre-trained model on eight tasks of the GLUE benchmark. We set the rank of the low-rank decomposition matrices to $4\sim 12$ and fine-tuned $20\%\sim 70\%$ of the modules in each layer of the pre-trained model. The experimental results are summarized in Table 8. It can be shown that our DiffoRA achieves the best results in almost all datasets. For instance, our approach outperforms the state-of-the-art method PAdapter on the CoLA dataset by 0.98%. Although DiffoRA achieves the second-best result on a few tasks, we achieve the highest average accuracy among all the baselines, i.e., 86.81% on average. The experimental results demonstrate the effectiveness of our method.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Our contributions and scope can be accurately reflected in the abstract and introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations can be found in Section 7.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We provide the full set of assumptions and a complete proof in the Appendix A.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provided a README file in our code for the reviewers to check.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provided an anonymous URL link for our codes in Section 5.1. All the datasets we used are public.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We have described our experimental settings in Section 5.1. For additional experimental settings on various benchmarks, we provided the details in Appendix D.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We have clearly stated the factors of variability that the error bars are capturing. We have run our experiments multiple times with different seeds and provided the mean and standard deviation values in Section 5.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We have provided the hardware configurations in Appendix D.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We confirm that we preserve the anonymity and conform to the NeurIPS Codes of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: This work is fundamental research in computer science and has no societal impact.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper poses no risks above.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All the codes and datasets used in this paper are open-sourced. We have properly cited these resources in our paper.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We have provided an anonymous URL link for our codes.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing or research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve crowdsourcing or research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We utilize LLM only for grammar and formatting purposes.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.